### v3

# A family of log-correlated Gaussian processes

Abstract.
A family of log-correlated Gaussian processes indexed by metric spaces is introduced, when the metric is conditionally negative definite. These processes arise as the limit of bi-fractional Brownian motions indexed by $(H,K)$ scaled by $K^{-1/2}$ as $K\downarrow 0$ with $H\in(0,1/2]$ fixed. When the metric is in addition a measure definite kernel, stochastic-integral representations of the generalized processes when evaluated at a test function are provided. The introduced processes are also shown to be the scaling limits of certain aggregated models.
1. Overview
Log-correlated Gaussian processes have received extensive attentions recently in probability theory and related fields. Such processes defined on a locally compact metric space $({\mathsf{M}},{\mathsf{d}})$ are generalized processes with covariance kernel in the form of
$\Gamma(x,y)=\log\frac{C}{{\mathsf{d}}(x,y)}+g(x,y),\quad x,y\in{\mathsf{M}},$
where $g(x,y)$ is a continuous function on ${\mathsf{M}}\times{\mathsf{M}}$. The corresponding Gaussian process is not point-wise defined since $\Gamma(x,x)=\infty$. In the seminal work, Kahane, [26] established a general theory of building the Gaussian multiplicative chaos starting from such kernels. Many developments have followed recently. See the survey of Rhodes and Vargas, [42] and the forthcoming monograph by Berestycki and Powell, [3]. In particular, many efforts have been devoted to specific log-correlated Gaussian processes, among which the most prominent example has been the Gaussian free fields [3, 46]. At the same time, log-correlated Gaussian processes and Gaussian multiplicative chaos have been shown to arise from limit theorems in random matrix theory [20, 6, 50] and mathematical finance [2, 37, 13]. Most of the time, $\mathsf{M}$} is either a domain of $\mathbb{R}^{n}$, or the circle ${\mathbb{S}}^{1}$. See also the survey by Duplantier et al., [8].
The goal of this paper is to introduce a family of log-correlated Gaussian processes indexed by metric spaces with a covariance function explicitly determined by the metric $\mathsf{d}$} and a parameter $H\in(0,1/2]$. For Brownian motion, Paul Lévy already investigated the question of defining Brownian motion indexed by other spaces than $\mathbb{R}$ [31, 36, 16]. Later on, the question of defining fractional Brownian motions indexed by other spaces including Euclidean spaces $\mathbb{R}^{d}$, Euclidean spheres ${\mathbb{S}}^{d}$, and hyperbolic spaces $\mathbb{H}^{d}$ with Hurst index $H\in(0,1/2)$ has been addressed, and these processes indexed by metric spaces can be characterized as centered Gaussian processes $G^{H}$ with ${\mathbb{E}}(G^{H}(x)-G^{H}(y))^{2}={\mathsf{d}}^{2H}(x,y)$ and $G^{H}(o)=0$ for some $o\in{\mathsf{M}}$ [24, 23, 25, 7]. A closely related question is the existence of the so-called bi-fractional Brownian motions indexed by metric spaces. A bi-fractional Brownian motion is a Gaussian process with two parameters that we let $H,K$ denote throughout. We shall see that for a large family of examples with $H\in(0,1/2]$, a log-correlated Gaussian process arises as the simple limit of bi-fractional Brownian motions with parameters $(H,K)$ normalized by $K^{-1/2}$ as $K\downarrow 0$. This, actually, follows from a simple calculation that we now explain.
1.1. From the bi-fractional Brownian motion to a log-correlated Gaussian process
We let $G^{H,K}=(G^{H,K}(x))_{x\in{\mathsf{M}}}$ denote the bi-fractional Brownian motion indexed by a metric space $\mathsf{M}$}, and $H>0,K>0$ are two parameters that we shall discuss in a moment. This is a centered Gaussian process with covariance function
(1.1) ${\rm{Cov}}\left(G^{H,K}(x),G^{H,K}(y)\right)=\frac{1}{2^{K}}\left(\left({\mathsf{d}}^{2H}(o,x)+{\mathsf{d}}^{2H}(o,y)\right)^{K}-{\mathsf{d}}^{2HK}(x,y)\right),x,y\in{\mathsf{M}},$
where $o\in{\mathsf{M}}$ is fixed and plays the role of the origin. Bi-fractional Brownian motion was first introduced by Houdré and Villa, [19] with $\mathsf{M}$$\mathbb{R}$ (and in this case $G^{H,K}$ is well-defined with $H\in(0,1],K\in(0,1]$ and $o=0$). In particular, $G^{H,1}$ is just a fractional Brownian motion with Hurst index $H\in(0,1]$ ($H=1$ is a degenerate case). The exact range of legitimate $(H,K)$ for the existence of $G^{H,K}$ remains an open question [49] even when $\mathsf{M}$$\mathbb{R}$ and $\mathsf{d}$} is the Euclidean metric; see also Lifshits and Volkova, [32].
Since then, bi-fractional Brownian motions indexed by various metric spaces have been studied; see Ma, [33]. In particular, when setting $K=1$ one always recover a fractional Brownian motion indexed by the same metric space with Hurst index $H$. It is important to keep in mind that the legitimate choice of $(H,K)$ depends on $({\mathsf{M}},{\mathsf{d}})$. As a general setup we assume that $\mathsf{d}$} is conditionally negative definite (a.k.a. of negative type), and in this case it is easy to check that the function on the right-hand side of (1.1) is positive definite and hence $G^{H,K}$ is well-defined with
$H\in(0,1/2]\quad\mbox{ and }\quad K\in(0,1].$
See Ma, [33, Theorem 1.1, (vii)]. In particular, $G^{H,1}$ is a fractional Brownian motion with Hurst index $H$ and $G^{1/2,1}$ is a Brownian motion.
Then, with $H$ fixed we have
(1.2) $\lim_{K\downarrow 0}\frac{1}{K}{\rm{Cov}}\left(G^{H,K}(x,y)\right)=\log\frac{{\mathsf{d}}^{2H}(o,x)+{\mathsf{d}}^{2H}(o,y)}{{\mathsf{d}}^{2H}(x,y)}=:\Gamma^{(H)}(x,y),x,y\in{\mathsf{M}}.$
This simple calculation deserves to be stated in a proposition.
Proposition 1.1.
If $({\mathsf{M}},{\mathsf{d}})$ is a metric space and $\mathsf{d}$} is of negative type, then for all $H\in(0,1/2]$, $\Gamma^{(H)}$ in (1.2) is a covariance kernel for a generalized Gaussian process indexed by $\mathsf{M}$}.
We let $G^{H}$ denote the so-introduced Gaussian process, and name it as the log-correlated bi-fractional Brownian motion with parameter $H\in(0,1/2]$.
Our first main contribution is to provide a stochastic-integral representation of $G^{H}$. Strictly speaking, we represent $G^{H}(f)$ (when evaluated at a suitable test function $f:{\mathsf{M}}\to\mathbb{R}$) in the form of $G^{H}(f)=\int_{\mathsf{M}}f(x)M({\rm d}x)$ where $M$ is a suitable Gaussian random measure. For this purpose, we shall impose a further condition on $\mathsf{d}$} being of negative type.
It has been recently shown that when $\mathsf{d}$} is a measure definite kernel, a notion from Robertson and Steger, [43], not only a fractional Brownian motion indexed by $\mathsf{M}$} with Hurst index $H\in(0,1/2)$ can be defined, but also it has a stochastic-integral representation of Chentsov-type [45]. Our results are closely related to this unified framework for fractional Brownian motions indexed by metric spaces [14]. The metric $\mathsf{d}$} being a measure definite kernel is also our key assumption here.
Assumption 1.2.
The metric $d$ on $\mathsf{M}$} is a measure definite kernel, in the sense that there exists a measure space $(E,{\mathcal{E}},\mu)$ and a family of sets $\{A_{x}\}_{x\in{\mathsf{M}}}\subset{\mathcal{E}}$ such that $\mu(A_{x})<\infty$ for all $x\in{\mathsf{M}}$ and
(1.3) ${\mathsf{d}}(x,y)=\mu(A_{x}\Delta A_{y}),\mbox{ for all }x,y\in{\mathsf{M}},$
where $A\Delta B=(A\setminus B)\cup(B\setminus A)$.
Recall that if $\mathsf{d}$} is of negative type then there exists a Gaussian process $(G(x))_{x\in{\mathsf{M}}}$ such that ${\mathbb{E}}(G(x)-G(y))^{2}={\mathsf{d}}(x,y)$. Every measure definite kernel is of negative type, but the opposite is not true. It has been well-known that if $\mathsf{d}$} is of negative type, then so is ${\mathsf{d}}^{\beta}$ for all $\beta\in(0,1)$. A brief review on metrics of negative type, measure definite kernels and their induced Gaussian processes can be found in Istas, [25]. It has been recently shown that if $\mathsf{d}$} is a measure definite kernel, then so is ${\mathsf{d}}^{\beta}$ for all $\beta\in(0,1)$ [14].
The first main result is the following.
Theorem 1.3.
If Assumption 1.2 is satisfied, then for $H\in(0,1/2]$,
$\Gamma^{(H)}(x,y)=\log\frac{\mu^{2H}(A_{x})+\mu^{2H}(A_{y})}{{\mathsf{d}}^{2H}(x,y)},x,y\in{\mathsf{M}},\beta\in(0,1],$
is the covariance kernel of a generalized Gaussian process on $\mathsf{M}$}. More precisely, we have
(1.4) $\Gamma^{(H)}(x,y)=\int_{0}^{\infty}\Gamma_{r}^{(H)}(x,y){\rm d}r$
where for each $r>0$, $\Gamma_{r}^{(H)}:{\mathsf{M}}\times{\mathsf{M}}\to\mathbb{R}$ is a positive definite function.
We shall provide an explicit expression for $\Gamma_{r}^{(H)}$, which is essentially based on a stochastic-integral representation of $G(f)$, in Section 4, Theorem 4.2.
Remark 1.4.
Note that $\Gamma^{(H)}$ in (1.4) is a $\sigma$-positive kernel in the sense of Kahane, [26], a property that is not easy to check in practice. The $\sigma$-positivity is a consequence of the metric $\mathsf{d}$} being a measure definite kernel (see (4.3)). Thanks to this property, the existence of the Gaussian multiplicative chaos determined by $\Gamma$ follows immediately from Kahane’s framework. We leave the investigations of so-obtained Gaussian multiplicative chaos for future studies.
Below we recall a few examples satisfying Assumption 1.2. If there is a point $o\in{\mathsf{M}}$ such that $\mu(A_{o})=0$, then the kernel simplifies to $\Gamma^{(H)}$ in (1.2). We also provide an example that satisfies Assumption 1.2 but cannot be obtained as the limit of bi-fractional Brownian motions in Example 1.3 below.
Example 1.1.
In the case $\mathsf{M}$$\mathbb{R}$, we take $A_{x}=[0,x]$ if $x\geq 0$ and (x,0](-x,0] if $x<0$, and $\mu$ as the Lebesgue measure. Then, we have
(1.5) $\Gamma^{(H)}(x,y)=\log\frac{|x|^{2H}+|y|^{2H}}{|x-y|^{2H}},x,y\in\mathbb{R}.$
Obviously, when $H=1/2$, $\Gamma(x,y)=0$ if $x<0<y$, which means that the restrictions of the process to $(-\infty,0)$ and $(0,\infty)$ respectively are independent. In Sections 2 and 3 we shall focus on this example with $H=1/2$ and restricted to $0,\infty)$.
Example 1.2.
For ${\mathsf{M}}=\mathbb{R}^{n}$, it is well-known that the Euclidean metric $\mathsf{d}$} is a measure-definite kernel. Indeed, one can take $E:=\{(s,r):s\in{\mathbb{S}}^{n},r\in(0,\infty)\}$ representing the space of all hyperplanes of $\mathbb{R}^{n}$ not including the origin, $A_{x}=\{(s,r):s\in{\mathbb{S}}^{n},0<r<\langle s,x\rangle\}$ the collection of all hyperplanes separating 0 and $x$ in $\mathbb{R}^{n}$, and $\mu({\rm d}s{\rm d}r)={\rm d}s{\rm d}r$ (${\rm d}s$ is the Lebesgue measure on ${\mathbb{S}}^{n}$, ${\rm d}r$ is the Lebesgue measure on $\mathbb{R}^{1}$, such that (1.3) holds. In fact, this representation of the measure definite kernel is at the heart of earlier Chentsov representation for fractional Lévy Brownian fields [45, Chapter 8.3]. So the covariance kernel is
$\Gamma^{(H)}(x,y)=\log\frac{{\left\|x\right\|}^{2H}+{\left\|y\right\|}^{2H}}{{\left\|x-y\right\|}^{2H}},\quad x,y\in\mathbb{R}^{n}.$
Example 1.3.
In the case ${\mathsf{M}}={\mathbb{S}}^{n}$ and $\mathsf{d}$} is the geodesic, we can take ${\mathcal{H}}_{x}$ as the hemisphere of ${\mathbb{S}}^{n}$ centered as $x\in{\mathbb{S}}^{n}$, $o\in{\mathbb{S}}^{n}$ a fixed point, $A_{x}:={\mathcal{H}}_{x}\Delta{\mathcal{H}}_{o},x\in{\mathbb{S}}^{n}$ and $\mu({\rm d}x)=\pi{\rm d}x/w_{n}$ where $w_{n}$ is the surface area of ${\mathbb{S}}^{n}$. In this way, (1.3) and (1.2) follow.
The above choice is well-known when studying the spherical fractional Brownian motion. Note that the spherical fractional Brownian motion and the corresponding log-correlated Gaussian process are not rotation-invariant. For a rotation-invariant model, one can consider simply $A_{x}={\mathcal{H}}_{x}$. This time, we still have ${\mathsf{d}}(x,y)=\mu({\mathcal{H}}_{x}\Delta{\mathcal{H}}_{y})$ (since $({\mathcal{H}}_{x}\Delta{\mathcal{H}}_{o})\Delta({\mathcal{H}}_{y}\Delta{\mathcal{H}}_{o})={\mathcal{H}}_{x}\Delta{\mathcal{H}}_{y}$). But now we have $\mu({\mathcal{H}}_{x})=\pi/2$ (compared with $\mu(A_{x})={\mathsf{d}}(o,x)$) and hence the corresponding generalized Gaussian process has covariance kernel
$\Gamma^{(H)}(x,y)=\log\frac{2(\pi/2)^{2H}}{{\mathsf{d}}^{2H}(x,y)}.$
For other examples, see [25, Section 2.8].
The fact that $G^{H}(f)$ has a stochastic-integral representation with $f$ from a suitable family of test functions yields immediately that the generalized Gaussian process $G$ can be extended to generalized infinitely-divisible processes [44] by replacing in the stochastic-integral representation the Gaussian random measure by infinitely-divisible ones (see [2, 41] for a closely related example in a similar spirit). We do not pursue the full generality in this direction here. Instead, we explain that the extension to stable ones is almost free. Last but not least, as the second main contribution we provide limit theorems for the introduced log-correlated bi-fractional Brownian motions.
The paper is organized as follows. A unified treatment for general $({\mathsf{M}},{\mathsf{d}})$ satisfying Assumption 1.2 and with $H\in(0,1/2]$ is provided in Section 4. However, the presentation might be a little too abstract for the first time reading. To help illustrate, we shall focus on Example 1.1, the log-correlated bi-fractional Brownian motion on $\mathbb{R}_{+}=[0,\infty)$ with covariance function (1.5) and $H=1/2$. For this case, in Section 2 we focus on stochastic-integral representations, and in Section 3 as the second main contribution we provide a central limit theorem based on a variation of a model by Enriquez, [11] when studying the fractional Brownian motion. This is also the case closely related to a few very recent developments in the literature; see Remarks 1.5 and 1.6. In Section 4, after presenting the results in the general framework, we also provide a central limit theorem for $G$ based on another model similar to the one in Fu and Wang, [14].
1.2. Related results in the literature
We conclude the introduction with a few remarks.
Remark 1.5.
In an inspiring paper, Neuman and Rosenbaum, [37] considered to characterize the limit process of fractional Brownian motions, denoted by ${\mathbb{B}}^{H}$, as $H\downarrow 0$. Therein, the authors investigated
(1.6) $X_{H}(t):=\frac{1}{\sqrt{H}}\left({\mathbb{B}}^{H}_{t}-\frac{1}{t}\int_{0}^{t}{\mathbb{B}}_{s}^{H}{\rm d}s\right),$
provided an argument on why a shift also makes sense in practice in finance applications, and then showed the limit of $X_{H}$ is a log-correlated Gaussian process. Later on, Hager and Neuman, [18] showed that the random shift $t^{-1}\int_{0}^{t}{\mathbb{B}}_{s}^{H}{\rm d}s$ can be relaced by a larger family of processes and the limits remain to be log-correlated (and also they considered $\mathbb{R}^{d}$-indexed random fields for $d\in\mathbb{N}$). Our process corresponds to a specific choice of the shift, essentially related to the decomposition of a fractional Brownian motion by a bi-fractional Brownian motion and a smooth Gaussian process due to Lei and Nualart, [30]. We provide more details in Remark 2.6.
Remark 1.6.
The limit process $G^{1/2}$ introduced here can already be read from recent developments by Iksanov and Kotelnikova, [22] on the Karlin model [28, 17], and it is from their results we first had the guess of the expression of $G^{1/2}$. The Karlin model is an infinite urn model of which the law of the sampling over $\mathbb{N}$ decays as $j^{-1/{2H}},H\in(0,1/2)$ as $j\to\infty$. The model was introduced by Karlin, [28], who showed that the partial sum of so-called odd-occupancy process (the counting process of urns occupied by an odd number of balls), normalized by $n^{H}$, scales to a Gaussian random variable. The consideration of odd-occupancy process was inspired by Spitzer, [48] to modeling the on/off events for lightbulbs. It is known that randomized odd-occupancy processes scale to a fractional Brownian motion with Hurst index $H\in(0,1/2)$ [10], and this type of limit theorems has been extended to stable processes indexed general metric spaces [9, 14]. The odd-occupancy mechanism is reflected in the stochastic-integral representations (see (2.16) below), and also in our representation for log-correlated bi-fractional Brownian motions.
The Karlin model with $H=0$ has been recently investigated by Iksanov et al., [21] and Iksanov and Kotelnikova, [22]. Limit theorems have been established in this regime with new Gaussian processes in the limit in the aforementioned papers. Their representation is not immediately recognizable to be related to ours and we provide some details in Appendix A. It should be expected that from their model a counterpart of our Theorem 3.1 can be established (they examined $j$-occupancy processes instead of odd-occupancy processes).
Remark 1.7.
Forde et al., [13] considered the limit of re-scaled Riemann–Liouville process and showed that the limit is a generalized Gaussian process with covariance function
$\log{\frac{\sqrt{t}+\sqrt{s}}{\sqrt{t}-\sqrt{s}}},\quad 0\leq s<t<\infty.$
Their limit theorem is in the same spirit as Neuman and Rosenbaum, [37].
Lifshits and Volkova, [32] also considered a limit theorem for $G^{H,K}$ as $K\downarrow 0$, although a Lamperti’s transformation is first applied to $G^{H,K}$ so that it is transformed to a stationary process before the limit is taken. A generalized process again arises in the limit.
Another example of a fractional Brownian motion with Hurst index $H=0$ has been seen in the studies of random matrix theory [15, 29]. The corresponding example is a centered self-similar Gaussian process with stationary increments.
Acknowledgements
Y.W. is grateful to Mikhail Lifshits for several stimulating discussions on an earlier version of the paper, which lead to an enhancement of the results and a major revision of the paper. Y.W. thanks also Alexander Iksanov, Joseph Najnudel, and Yi Shen for several discussions. Y.W. was partially supported by Simons Foundation (MP-TSM-00002359), and a Taft Center Fellowship (2024–2025) from Taft Research Center at University of Cincinnati.
2. A stochastic-integral representation
In this and the next sections we consider ${\mathsf{M}}=[0,\infty)$, ${\mathsf{d}}(x,y)=|x-y|$. In most of this section we focus $H=1/2$ in the parameter of the bi-fractional Brownian motion except in Section 2.2.
Our reference for stochastic integrals with respect to stable random measures is Samorodnitsky and Taqqu, [45]: if $(E,{\mathcal{E}},\mu)$ is a measure space, $M_{\alpha}$ is a symmetric $\alpha$-stable random measure on $(E,{\mathcal{E}})$ with control measure $\mu$, then the stochastic integral $\int f{\rm d}M_{\alpha}$ with control measure $\mu$ is defined as a symmetric $\alpha$-stable random variable with
(2.1) ${\mathbb{E}}\exp\left({\rm i}\theta\int f{\rm d}M_{\alpha}\right)=\exp\left(-|\theta|^{\alpha}\int|f|^{\alpha}{\rm d}\mu\right),\mbox{ for all }f\in L^{\alpha}(E,\mu),\theta\in\mathbb{R}.$
We first focus on $\alpha=2$, and in this case $\int f{\rm d}M_{2}$ is a centered Gaussian random variable and ${\rm{Cov}}(\int f{\rm d}M_{2},\int g{\rm d}M_{2})=2\int fg{\rm d}\mu$ for $f,g\in L^{2}(E,\mu)$. (Note that the multiplicative constant $2$ is introduced to fit into the unified formula in (2.1).)
Let $(\Omega,{\mathcal{F}},\mathbb{P})$ be the by-default probability space that the random measures are defined on. Set $E=\mathbb{R}_{+}\times\Omega^{\prime}$ where $(\Omega^{\prime},{\mathcal{F}}^{\prime},\mathbb{P}^{\prime})$ is another probability space on which a standard Poisson process $N^{\prime}$ is defined, and $M_{2}$ is a Gaussian random measure on $\mathbb{R}_{+}\times\Omega^{\prime}$ with control measure $2r^{-1}{\rm d}r{\rm d}\mathbb{P}^{\prime}$. Write
$\widetilde{p}_{r}(t)=\mathbb{P}^{\prime}(N^{\prime}(rt){\rm~is~odd})=\frac{1}{2}\left(1-e^{-2rt}\right).$
We introduce
(2.2) $G(f):=\int_{\mathbb{R}_{+}\times\Omega^{\prime}}\int_{\mathbb{R}_{+}}\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\widetilde{p}_{r}(t)\right)f(t){\rm d}tM_{2}({\rm d}r,{\rm d}\omega^{\prime}),$
and
(2.3) ${\mathcal{F}}_{0}:=\left\{f\in L^{\infty}([0,\infty)):\int_{1}^{\infty}|f(s)|\log s{\rm d}s<\infty\right\}.$
Lemma 2.1.
The collection $\{G(f)\}_{f\in{\mathcal{F}}_{0}}$ is a family of centered Gaussian random variables with covariance function
${\rm{Cov}}(G(f),G(g))=\int_{0}^{\infty}\int_{0}^{\infty}f(s)g(t)\log\frac{s+t}{|s-t|}{\rm d}s{\rm d}t,\quad\mbox{ for all }f,g\in{\mathcal{F}}_{0},$
that also satisfies
(2.4) $G(a_{1}f_{1}+a_{2}f_{2})=a_{1}G(f_{1})+a_{2}G(f_{2})\mbox{ almost surely, for all }a_{1},a_{2}\in\mathbb{R},f_{1},f_{2}\in{\mathcal{F}}_{0}.$
Proof.
We claim that
${\rm{Cov}}(G(f),G(f))$ $=2\int_{\mathbb{R}_{+}\times\Omega^{\prime}}\left(\int_{0}^{\infty}\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\widetilde{p}_{r}(t)\right)f(t){\rm d}t\right)^{2}\frac{2{\rm d}r}{r}{\rm d}\mathbb{P}^{\prime}$
(2.5) $=\int_{\mathbb{R}_{+}^{2}}f(s)f(t)\Gamma(s,t){\rm d}s{\rm d}t.$
The first equality follows from the property of stochastic integrals. Next we show the expression on the right-hand side of (2.5) is finite. We prove $\int_{0}^{\infty}\int_{s}^{\infty}|f(s)f(t)\log((s+t)/(t-s))|{\rm d}t{\rm d}s<\infty$. To see this, we observe
$\int_{0}^{\infty}\int_{s+1}^{\infty}|f(s)f(t)|\left|\log{\frac{s+t}{t-s}}\right|{\rm d}t{\rm d}s\leq\int_{0}^{\infty}|f(s)|\int_{s+1}^{\infty}|f(t)|\log(4t){\rm d}t{\rm d}s<\infty,$
$\int_{0}^{\infty}\int_{s}^{s+1}\left|f(s)f(t)\log\frac{1}{t-s}\right|{\rm d}t{\rm d}s\leq{\left\|f\right\|}_{\infty}\int_{0}^{\infty}|f(s)|{\rm d}s\int_{0}^{1}|\log t|{\rm d}t<\infty,$
$\int_{1}^{\infty}\int_{s}^{s+1}\left|f(s)f(t)\right|\log(s+t){\rm d}t{\rm d}s\leq\int_{1}^{\infty}\int_{s}^{s+1}|f(s)f(t)|\log(2t){\rm d}t{\rm d}s<\infty,$
and
$\int_{0}^{1}\int_{s}^{s+1}\left|f(s)f(t)\log(s+t)\right|{\rm d}t{\rm d}s$ $\leq{\left\|f\right\|}_{\infty}^{2}\int_{0}^{1}\int_{s}^{s+1}\left|\log(s+t)\right|{\rm d}t{\rm d}s$
$\leq{\left\|f\right\|}_{\infty}^{2}\int_{0}^{1}|\log s|+\log(1+s){\rm d}s<\infty.$
It remains to prove the equality in (2.5), which follows from a careful application of Fubini’s theorem. Indeed, for all $0\leq s<t$,
(2.6) ${\mathbb{E}}^{\prime}\left(\left({{\bf 1}}_{\left\{N^{\prime}(rs){\rm~is~odd}\right\}}-\widetilde{p}_{s}(r)\right)\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\widetilde{p}_{t}(r)\right)\right)\\ =\frac{1}{4}\left((1-e^{-2rs})(1+e^{-2r(t-s)})-(1-e^{-2rs})(1-e^{-2rt})\right)=\frac{1}{4}\left(e^{-2r(t-s)}-e^{-2r(t+s)}\right).$
We then start by rewriting the expression after first equality in (2.5) as
$4\int_{\mathbb{R}_{+}}$ $\int_{\Omega^{\prime}}\int_{\mathbb{R}_{+}^{2}}\left({{\bf 1}}_{\left\{N^{\prime}(rs){\rm~is~odd}\right\}}-\widetilde{p}_{r}(s)\right)\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\widetilde{p}_{r}(t)\right)f(s)f(t){\rm d}s{\rm d}t{\rm d}\mathbb{P}^{\prime}\frac{{\rm d}r}{r}$
$=4\int_{\mathbb{R}_{+}}\int_{\mathbb{R}_{+}^{2}}\int_{\Omega^{\prime}}\left({{\bf 1}}_{\left\{N^{\prime}(rs){\rm~is~odd}\right\}}-\widetilde{p}_{r}(s)\right)\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\widetilde{p}_{r}(t)\right)f(s)f(t){\rm d}\mathbb{P}^{\prime}{\rm d}s{\rm d}t\frac{{\rm d}r}{r}$
$=\int_{\mathbb{R}_{+}}\int_{\mathbb{R}_{+}^{2}}\frac{e^{-2r|t-s|}-e^{-2r(t+s)}}{r}f(s)f(t){\rm d}s{\rm d}t{\rm d}r.$
Here in the first step we applied Fubini’s theorem to interchange the two inner integrals (in this step when verifying the condition in Fubini’s theorem for each $r>0$, we bounded $|{{\bf 1}}_{\left\{N^{\prime}(rs){\rm~is~odd}\right\}}-\widetilde{p}_{r}(s)||{{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\widetilde{p}_{r}(t)|\leq 1$ and used the assumption ${\left\|f\right\|}_{1}<\infty$), and in the second step we applied (2.6). In the last expression above, we can interchange the two integrals again by Fubini’s theorem, which, in combination of Frullani integral
(2.7) $\int_{0}^{\infty}\frac{e^{-r(t-s)}-e^{-r(t+s)}}{r}{\rm d}r=\log{\frac{t+s}{t-s}},\quad 0\leq s<t,$
leads to the equality in (2.5) (in this step we also need the fact that the right-hand side of (2.5) is finite, which we have shown earlier).
Once we have shown that (2.2) is a well-defined stochastic integral, the linearity relation (2.4) holds immediately by linearity of stochastic integrals. Then using ${\rm{Cov}}(X,Y)=({\rm{Cov}}(X+Y,X+Y)-{\rm{Cov}}(X,X)-{\rm{Cov}}(Y,Y))/2$ we obtain the stated covariance formula. ∎
Next we introduce an approximation of $G(f)$. For each $\epsilon>0$, set
(2.8) $G^{(\epsilon)}(f):=\int_{\mathbb{R}_{+}\times\Omega^{\prime}}\int_{\mathbb{R}_{+}}\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\widetilde{p}_{r}(t)\right){{\bf 1}}_{\left\{r\leq\epsilon^{-1}\right\}}f(t){\rm d}tM_{2}({\rm d}r{\rm d}\omega^{\prime}),$
where we follow the same notations for stochastic integrals earlier. Note that by stochastic Fubini theorem [38, Proposition 5.13.1], we have the representation
(2.9) $G^{(\epsilon)}(f)=\int_{0}^{\infty}\widetilde{G}^{(\epsilon)}_{t}f(t){\rm d}t,\mbox{ almost surely,}$
where
$\widetilde{G}^{(\epsilon)}_{t}:=\int_{\mathbb{R}_{+}\times\Omega^{\prime}}\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\mathbb{P}^{\prime}(N^{\prime}(rt){\rm~is~odd})\right){{\bf 1}}_{\left\{r\leq\epsilon^{-1}\right\}}M_{2}({\rm d}r{\rm d}\omega^{\prime}),t\geq 0,$
is a point-wise defined stochastic process. The identity (2.9) is often referred in the literature to as the point-wise representation of $G^{(\epsilon)}$ as a generalized function.
Lemma 2.2.
  1. (i)
    For each $\epsilon>0$, the process $\{\widetilde{G}_{t}^{(\epsilon)}\}_{t\geq 0}$ is a centered Gaussian process with
    ${\rm{Cov}}\left(\widetilde{G}^{(\epsilon)}_{s},\widetilde{G}^{(\epsilon)}_{t}\right)=\int_{0}^{\epsilon^{-1}}\frac{e^{-2r|t-s|}-e^{-2r(t+s)}}{r}{\rm d}r.$
  2. (ii)
    For each $\epsilon>0$, we have
    ${\rm{Cov}}\left(G^{(\epsilon)}(f),G^{(\epsilon)}(g)\right)=\int_{0}^{\infty}\int_{0}^{\infty}f(s)g(t)\int_{0}^{\epsilon^{-1}}\frac{e^{-2r|t-s|}-e^{-2r(t+s)}}{r}{\rm d}r{\rm d}s{\rm d}t,f,g\in{\mathcal{F}}_{0}.$
    In particular, $\int_{0}^{\infty}\widetilde{G}_{t}^{(\epsilon)}f(t){\rm d}t\Rightarrow G(f)$ as $\epsilon\downarrow 0$ for all $f\in{\mathcal{F}}_{0}$.
Proof.
For the first part, we have
${\rm{Cov}}\left(\widetilde{G}^{(\epsilon)}_{s},\widetilde{G}^{(\epsilon)}_{t}\right)$ $=4\int_{[0,\epsilon^{-1}]\times\Omega^{\prime}}\left({{\bf 1}}_{\left\{N^{\prime}(rs){\rm~is~odd}\right\}}-\widetilde{p}_{s}(r)\right)\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\widetilde{p}_{t}(r)\right)\frac{{\rm d}r}{r}\mathbb{P}^{\prime}({\rm d}\omega^{\prime})$
(2.10) $=4\int_{0}^{\epsilon^{-1}}{\mathbb{E}}^{\prime}\left(\left({{\bf 1}}_{\left\{N^{\prime}(rs){\rm~is~odd}\right\}}-\widetilde{p}_{s}(r)\right)\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\widetilde{p}_{t}(r)\right)\right)\frac{{\rm d}r}{r},$
where in (2.10) we applied Fubini’s theorem. In fact, by a straightforward calculation we have
${\mathbb{E}}\left|\left({{\bf 1}}_{\left\{N^{\prime}(rs){\rm~is~odd}\right\}}-\widetilde{p}_{s}(r)\right)\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\widetilde{p}_{t}(r)\right)\right|=\frac{1}{4}\left(1-e^{-4rs}\right).$
(This function is not integrable with respect to ${\rm d}r/r$ at infinity, and hence we introduced the truncation ${{\bf 1}}_{\left\{r\leq\epsilon^{-1}\right\}}$.) The covariance function in the second part follows from a similar calculation. The convergence in fact can be read directly from (2.8) by letting $\epsilon\downarrow 0$ (compared with (2.2)). ∎
2.1. An extension to a generalized stable process
The stochastic-integral representation (2.2) provides an immediate extension to a generalized infinitely-divisible process [44] by replacing the Gaussian random measure accordingly. We only describe an extension to stable ones for the sake of simplicity. For each $\alpha\in(0,2)$, we extend the definition of $G(f)$ in (2.2) to
$G_{\alpha}(f):=\int_{\mathbb{R}_{+}\times\Omega^{\prime}}\int_{\mathbb{R}_{+}}\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\widetilde{p}_{r}(t)\right)f(t){\rm d}tM_{\alpha}({\rm d}r,{\rm d}\omega^{\prime}),$
where $M_{\alpha}$ is a symmetric $\alpha$-stable random measure on $0,\infty)\times\Omega^{\prime}$ with control measure $2r^{-1}{\rm d}r{\rm d}\mathbb{P}^{\prime}$. So $G(f)=G_{2}(f)$ and we again have $G_{\alpha}(a_{1}f_{1}+a_{2}f_{2})=a_{1}G_{\alpha}(f_{1})+a_{2}G_{\alpha}(f_{2})$ almost surely for $a_{1},a_{2}\in\mathbb{R},f_{1},f_{2}$ from a suitably chosen class of functions. Recall that $M_{\alpha}$ is a symmetric $\alpha$-stable random measure and the corresponding stochastic integrals are random variables characterized by (2.1). Formally, we have
(2.11) ${\mathbb{E}}\exp\left({\rm i}G_{\alpha}(f)\right)=\exp\left(-2\int_{0}^{\infty}{\mathbb{E}}^{\prime}\left|\int_{0}^{\infty}\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\widetilde{p}_{r}(t)\right)f(t){\rm d}t\right|^{\alpha}\frac{{\rm d}r}{r}\right).$
Note that when $\alpha\in(0,2)$ it is not immediately clear what is the exact class of functions so that the right-hand side above is finite. Recall ${\mathcal{F}}_{0}$ in (2.3) and introduce
(2.12) ${\mathcal{F}}_{\delta}:=\left\{f\in L^{\infty}([0,\infty)):\int_{1}^{\infty}s^{\delta}|f(s)|{\rm d}s<\infty\right\},\delta>0.$
In particular, $f\in{\mathcal{F}}_{\delta}$ for any $\delta\geq 0$ implies that ${\left\|f\right\|}_{1}<\infty$.
Lemma 2.3.
Assume one of the following:
  1. (i)
    $\alpha\in(0,2],f\in{\mathcal{F}}_{\delta}$ for some $\delta>0$, or
  2. (ii)
    $\alpha\in(1,2],f\in{\mathcal{F}}_{0}$.
Then, $G_{\alpha}(f)$ is a symmetric $\alpha$-stable random variable with characteristic function (2.11) (which is finite).
Proof.
Again, by properties of stochastic integrals, it is equivalent to show that
(2.13) $\int_{\mathbb{R}_{+}\times\Omega^{\prime}}\left|\int_{\mathbb{R}_{+}}\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\widetilde{p}_{r}(t)\right)f(t){\rm d}t\right|^{\alpha}\frac{{\rm d}r}{r}{\rm d}\mathbb{P}^{\prime}\\ =\int_{0}^{\infty}{\mathbb{E}}^{\prime}\left|\int_{0}^{\infty}\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\widetilde{p}_{r}(t)\right)f(t){\rm d}t\right|^{\alpha}\frac{{\rm d}r}{r}<\infty.$
Without loss of generality, assuming $f\geq 0$. Introduce
$\Psi_{f}(r):=\int_{0}^{\infty}\int_{s}^{\infty}f(s)f(t)\left(e^{-r(t-s)}-e^{-r(t+s)}\right){\rm d}t{\rm d}s.$
We shall show
(2.14) $\Psi_{f}(r)\leq\begin{cases}C(r^{\delta}\wedge r^{-1}),&\mbox{ if }f\in{\mathcal{F}}_{\delta},\delta>0,\\ C(|\log r|^{-2}\wedge r^{-1}),&\mbox{ if }f\in{\mathcal{F}}_{0},\end{cases}\quad\mbox{ for all }r>0,$
where the constant $C$ depends on $f$ but not $r$. Then, by Hölder’s inequality, we have
${\mathbb{E}}\left|\int_{0}^{\infty}\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\widetilde{p}_{r}(t)\right)f(t){\rm d}t\right|^{\alpha}$ $\leq\left({\mathbb{E}}\left|\int_{0}^{\infty}\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\widetilde{p}_{r}(t)\right)f(t){\rm d}t\right|^{2}\right)^{\alpha/2},$
$=\left(2\Psi_{f}(r)\right)^{\alpha/2},$
and hence (2.13) becomes $\int_{0}^{\infty}\Psi_{f}(r)^{\alpha/2}r^{-1}{\rm d}r<\infty$, which now follows from (2.14).
It remains to prove (2.14). First, for all $r>0$ we have
$\Psi_{f}(r)\leq\int_{0}^{\infty}f(s)\int_{0}^{\infty}f(s+t)e^{-rt}{\rm d}t{\rm d}s\leq{\left\|f\right\|}_{\infty}\frac{1}{r}\int_{0}^{\infty}f(s){\rm d}s\leq{\left\|f\right\|}_{\infty}{\left\|f\right\|}_{1}r^{-1}.$
Next, assuming $f\in{\mathcal{F}}_{\delta}$ with $\delta>0$ first. By a change of variables write first
$\Psi_{f}(r)=\int_{0}^{\infty}\int_{0}^{\infty}f(s)f(s+t)e^{-rt}(1-e^{-2rs}){\rm d}t{\rm d}s.$
Next we prove for all $r>0$, $\Psi_{f}(r)\leq(2{\left\|f\right\|}_{1}\int_{0}^{\infty}s^{\delta}f(s){\rm d}s)r^{\delta}$. Indeed, we have
$\int_{0}^{1/r}\int_{0}^{\infty}f(s)f(s+t)e^{-rt}(1-e^{-2rs}){\rm d}t{\rm d}s$ $\leq 2r\int_{0}^{1/r}sf(s)\int_{s}^{\infty}f(t){\rm d}t{\rm d}s$
$\leq 2r{\left\|f\right\|}_{1}\int_{0}^{1/r}(1/r)^{1-\delta}s^{\delta}f(s){\rm d}s$
$\leq 2r^{\delta}{\left\|f\right\|}_{1}\int_{0}^{\infty}s^{\delta}f(s){\rm d}s,$
and
$\int_{1/r}^{\infty}\int_{0}^{\infty}f(s)f(s+t)e^{-rt}(1-e^{-2rs}){\rm d}t{\rm d}s$ $\leq\int_{1/r}^{\infty}f(s)\int_{s}^{\infty}f(t){\rm d}t{\rm d}s$
$\leq{\left\|f\right\|}_{1}\int_{1/r}^{\infty}(1/r)^{-\delta}s^{\delta}f(s){\rm d}s$
$\leq r^{\delta}{\left\|f\right\|}_{1}\int_{0}^{\infty}s^{\delta}f(s){\rm d}s.$
We have proved (2.14) with $f\in{\mathcal{F}}_{\delta},\delta>0$.
Now assume $f\in{\mathcal{F}}_{0}$. This time, fix $\gamma\in(0,1)$. We have,
$\int_{0}^{(1/r)^{\gamma}}\int_{0}^{\infty}f(s)f(s+t)e^{-rt}(1-e^{-2rs}){\rm d}t{\rm d}s$ $\leq 2r\int_{0}^{(1/r)^{\gamma}}sf(s)\int_{s}^{\infty}f(t){\rm d}t{\rm d}s$
$\leq 2r^{1-\gamma}{\left\|f\right\|}_{1}^{2},$
and, for $r<1/2$,
$\int_{(1/r)^{\gamma}}^{\infty}f(s)\int_{s}^{\infty}f(t){\rm d}t{\rm d}s$ $\leq\int_{(1/r)^{\gamma}}^{\infty}\frac{f(s)}{\log s}\int_{s}^{\infty}f(t)\log t{\rm d}t{\rm d}s$
$\leq\frac{1}{\gamma^{2}\log^{2}r}\int_{1/r}^{\infty}f(s)\log s{\rm d}s\int_{s}^{\infty}f(t)\log t{\rm d}t\leq C\frac{1}{\log^{2}r}.$
So for $r<1/2$ we have $\Psi_{f}(r)<C/\log^{2}r$. By possibly increasing $C$ we have proved (2.14) when $f\in{\mathcal{F}}_{0}$. ∎
2.2. A representation for the bi-fractional Brownian motion with $H\in(0,1/2)$
In the previous discussions in this section, we essentially worked with $G^{1/2,2H}$, of which the stochastic-integral representations have been known. Here, we provide a new stochastic-integral representation for $G^{H,K}$ with $H\in(0,1/2)$ (to be compared with the case $H=1/2$, which we recall in (2.17) below). Consider
$\widetilde{G}^{H,K}(t)=\int_{\mathbb{R}_{+}\times\Omega^{\prime}}\left({{\bf 1}}_{\left\{N^{\prime}(\mathcal{S}_{2H,r}^{\prime}t){\rm~is~odd}\right\}}-\widetilde{p}_{2H,r}(t)\right)\widetilde{M}_{2,K}({\rm d}r{\rm d}\omega^{\prime}),$
where this time $\widetilde{M}_{2,K}$ is on $\mathbb{R}_{+}\times\Omega^{\prime}$ with control measure $C_{K}r^{-K-1}{\rm d}r{\rm d}\mathbb{P}^{\prime}$ where $\mathbb{P}^{\prime}$ is a probability measure on $\Omega^{\prime}$, under which $N^{\prime}$ is a standard Poisson process and $\mathcal{S}_{2H,r}$ is such that ${\mathbb{E}}^{\prime}e^{-\theta\mathcal{S}^{\prime}_{2H,r}}=e^{r\theta^{2H}}$ for all $\theta\geq 0$, $N^{\prime}$ and $\mathcal{S}_{2H,r}^{\prime}$ are independent, and
$\widetilde{p}_{2H,r}(t):=\mathbb{P}^{\prime}\left(N^{\prime}(\mathcal{S}_{2H,r}^{\prime}t){\rm~is~odd}\right)={\mathbb{E}}^{\prime}\left(\frac{1}{2}\left(1-e^{-2\mathcal{S}^{\prime}_{2H,r}t}\right)\right)=\frac{1}{2}\left(1-e^{-(2t)^{2H}r}\right).$
Lemma 2.4.
Under the notations above, for all $H\in(0,1/2]$ and $K\in(0,1)$ we have
${\rm{Cov}}\left(\widetilde{G}^{H,K}(s),\widetilde{G}^{H,K}(t)\right)=\frac{2^{(2H-1)K}}{4}\left((s^{2H}+t^{2H})^{K}-(t-s)^{2HK}\right),0\leq s<t<\infty.$
Proof.
Assume $s<t$. We first compute (with respect to $\mathbb{P}^{\prime}$)
${\rm{Cov}}^{\prime}\left({{\bf 1}}_{\left\{N^{\prime}(\mathcal{S}_{2H,r}^{\prime}s){\rm~is~odd}\right\}},{{\bf 1}}_{\left\{N^{\prime}(\mathcal{S}_{2H,r}^{\prime}t){\rm~is~odd}\right\}}\right)=\frac{1}{4}\left(e^{-2^{2H}(t-s)^{2H}r}-e^{-2^{2H}(s^{2H}+t^{2H})r}\right).$
Then,
${\rm{Cov}}\left(\widetilde{G}^{H,K}(s),\widetilde{G}^{H,K}(t)\right)$ $=C_{K}\int_{0}^{\infty}r^{-K-1}{\rm d}r{\rm{Cov}}^{\prime}\left({{\bf 1}}_{\left\{N^{\prime}(\mathcal{S}_{2H,r}^{\prime}s){\rm~is~odd}\right\}},{{\bf 1}}_{\left\{N^{\prime}(\mathcal{S}_{2H,r}^{\prime}t){\rm~is~odd}\right\}}\right)$
$=\frac{C_{K}}{4}\int_{0}^{\infty}\left(e^{-2^{2H}(t-s)^{2H}r}-e^{-2^{2H}(s^{2H}+t^{2H})r}\right)r^{-K-1}{\rm d}r$
$=\frac{C_{K}}{4}\frac{\Gamma(1-K)}{K}2^{2HK}\left(\left(s^{2H}+t^{2H}\right)^{K}-(t-s)^{2HK}\right).$
Note that when $H=1/2$, this representation was known [9, 14]. A new idea here for $H\in(0,1/2)$ is the introduction of a stable subordinator $\mathcal{S}_{2H,r}^{\prime}$ in the representation, and the same idea shall be applied in Section 4.
Remark 2.5.
In fact, one can have accordingly stochastic representations [14] for bi-fractional Brownian motions indexed by metric spaces. Furthermore, replacing Gaussian random measures by symmetric $\alpha$-stable ones, we obtain immediately extensions of bi-fractional Brownian motions, which we may name bi-fractional stable motions. We omit the details.
Remark 2.6.
Our result can be interpreted as a limit theorem for ${\mathbb{B}}^{H}$ as $H\downarrow 0$ with a different shift from the one considered by Neuman and Rosenbaum, [37] in (1.6). Our choice of the shift is in fact related to the decomposition of the fractional Brownian motion by a bi-fractional Brownian motion due to Lei and Nualart, [30]. Namely, ${\mathbb{B}}^{H}$ satisfies
(2.15) $\left({\mathbb{B}}^{H}_{t}\right)_{t\in[0,\infty)}\stackrel{{\scriptstyle d}}{{=}}\left(G^{1/2,2H}_{t}+W^{H}_{t}\right)_{t\in[0,\infty)},$
where $G^{1/2,2H}$ is a bi-fractional Brownian motion introduced earlier and $W^{H}$ satisfies
${\rm{Cov}}\left(W^{H}_{s},W^{H}_{t}\right)=\frac{1}{2}\left(s^{2H}+t^{2H}-(s+t)^{2H}\right),$
and the two processes are independent. Moreover, the decomposition (2.15) can hold in the almost sure sense. Let $(\Omega,{\mathcal{F}},\mathbb{P})$ be the by-default probability space that the random measures are defined on. Set $E=\mathbb{R}_{+}\times\Omega^{\prime}$ where $(\Omega^{\prime},{\mathcal{F}}^{\prime},\mathbb{P}^{\prime})$ is another probability space on which a standard Poisson process $N^{\prime}$ is defined, and $\widetilde{M}_{2,H}$ as a Gaussian random measure (defined on $(\Omega,{\mathcal{F}},\mathbb{P})$) on $E=\mathbb{R}_{+}\times\Omega^{\prime}$ with control measure $C_{2H}r^{-2H-1}{\rm d}r{\rm d}\mathbb{P}^{\prime}$ with $C_{\beta}=\beta 2^{-\beta}/\Gamma(1-\beta),\beta\in(0,1)$. Then, (2.15) holds with the following:
(2.16) ${\mathbb{B}}_{t}^{H}$ $:=\int_{\mathbb{R}_{+}\times\Omega^{\prime}}{{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}\widetilde{M}_{2,2H}({\rm d}r{\rm d}\omega^{\prime}),$
(2.17) $G^{1/2,2H}_{t}$ $:=\int_{\mathbb{R}_{+}\times\Omega^{\prime}}\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\widetilde{p}_{r}(t)\right)\widetilde{M}_{2,2H}({\rm d}r{\rm d}\omega^{\prime}),$
$W^{H}_{t}$ $:=\int_{\mathbb{R}_{+}\times\Omega^{\prime}}\widetilde{p}_{r}(t)\widetilde{M}_{2,2H}({\rm d}r{\rm d}\omega^{\prime}),$
with $\widetilde{p}_{r}(t)=\mathbb{P}^{\prime}(N^{\prime}(rt){\rm~is~odd})=\left(1-e^{-2rt}\right)/2$ [9, 14].
Now, one can write
$G^{1/2,2H}_{t}={\mathbb{B}}^{H}_{t}-W_{t}^{H}.$
That is, $G^{1/2,2H}$ is ${\mathbb{B}}^{H}$ with a random shift $W^{H}$. The convergence of $H^{-1/2}G^{1/2,2H}\Rightarrow G^{1/2}$ is now a special case of (1.2).
3. A central limit theorem
Another feature of the introduced process $G\equiv G^{1/2}$ is that it arises from a simple aggregated model. The aggregated model here is a variation of the one proposed by Enriquez, [11] when investigating approximations of a fractional Brownian motion. Connections between aggregated models and fractional Gaussian processes and random fields with long-range dependence [39, 44] are well-known [4, 5, 27, 35] (in particular, [4, 5] established limit theorems for generalized processes).
We follow the notations in Shen et al., [47], where the Enriquez model was revisited and extended. The model is an aggregated model of an increasing number, denoted by $m_{n}$, of i.i.d. layers indexed by $i\in\mathbb{N}$, and each layer consists of a sequence of stationary random variables denoted by $(\mathcal{X}_{i,j})_{j\in\mathbb{N}}$ that we now introduce. First, let $\mathcal{X}$ be a standard Gaussian random variable, $q$ a uniform random variable taking values from $(0,1)$, and $\{\eta_{j}^{(q)}\}_{j\in\mathbb{N}}$ conditionally i.i.d. Bernoulli random variables with parameter $q$. Assume that $\mathcal{X}$ is independent from $q$ and $\{\eta_{j}^{(q)}\}_{j\in\mathbb{N}}$. Let $(\mathcal{X}_{i},q_{i},\{\eta_{i,j}^{(q_{i})}\}_{j\in\mathbb{N}})_{i\in\mathbb{N}}$ denote i.i.d. copies of $(\mathcal{X},q,\{\eta_{j}^{(q)}\}_{j\in\mathbb{N}})$, and set
$\tau_{i,j}^{(q_{i})}:=\sum_{k=1}^{j}\eta_{i,k}^{(q_{i})},j\in\mathbb{N}.$
Then, the $j$-th element of $i$-th layer is defined as
$\mathcal{X}_{i,j}:=\frac{\mathcal{X}_{i}}{q^{1/2}}\left((-1)^{\tau_{i,j}^{(q_{i})}+1}\eta_{i,j}^{(q_{i})}-{\mathbb{E}}\left((-1)^{\tau_{i,j}^{(q_{i})}+1}\eta_{i,j}^{(q_{i})}\;\middle|\;q_{i}\right)\right).$
Now, the partial-sum process of the aggregated model is represented as
$\widehat{S}_{n,j}:=\frac{1}{\sqrt{m_{n}}}\sum_{i=1}^{m_{n}}\sum_{k=1}^{j}\mathcal{X}_{i,k}=\frac{1}{\sqrt{m_{n}}}\sum_{i=1}^{m_{n}}\frac{\mathcal{X}_{i}}{q_{i}^{1/2}}\left({{\bf 1}}_{\left\{\tau_{i,j}^{(q_{i})}{\rm~is~odd}\right\}}-\mathbb{P}\left(\tau_{i,j}^{(q_{i})}{\rm~is~odd}\;\middle|\;q_{i}\right)\right),j\in\mathbb{N}.$
In order to see that the proposed model exhibits log-correlated dependence structure, a quick calculation yields
(3.1) $\lim_{n\to\infty}{\rm{Cov}}\left(\widehat{S}_{n}(s),\widehat{S}_{n}(t)\right)=\frac{1}{4}\log{\frac{s+t}{s-t}},0<s<t.$
Indeed,
${\rm{Cov}}\left(\widehat{S}_{n}(s),\widehat{S}_{n}(t)\right)$ $=\int_{0}^{1}\frac{1}{r}\Bigg[\left(\frac{1}{2}\left(1-\left(1-2r\right)^{\left\lfloor ns\right\rfloor}\right)\frac{1}{2}\left(1+(1-2r)^{\left\lfloor n(t-s)\right\rfloor}\right)\right)$
$\quad\quad-\frac{1}{4}\left(1-(1-2r)^{\left\lfloor ns\right\rfloor}\right)\frac{1}{2}\left(1-(1-2r)^{\left\lfloor nt\right\rfloor}\right)\Bigg]{\rm d}r$
$\to\frac{1}{4}\int_{0}^{\infty}\frac{1}{r}\left(e^{-2(t-s)r}-e^{-2(t+s)r}\right){\rm d}r=\frac{1}{4}\log{\frac{t+s}{t-s}}.$
(Note that for binomial random variable $\tau_{n}^{(p)}$ with parameter $(n,p)$, $\mathbb{P}(\tau_{n}^{(p)}{\rm~is~odd})=(1/2)(1-(1-2p)^{n})$ and $\mathbb{P}(\tau_{n}^{(p)}{\rm~is~even})=(1/2)(1+(1-2p)^{n})$.) The step of convergence $\to$’ would require a little extra work. This is a standard analysis and details are omitted here; we do not make use of (3.1) in the sequel.
The convergence in (3.1) suggests that there is no convergence of point-wise defined stochastic processes, and instead one should search for convergence in the sense of generalized processes. For this purpose, recall ${\mathcal{F}}_{\delta}$ in (2.12) and introduce
(3.2) $G_{n}(f):=2\sum_{j=1}^{\infty}\widehat{S}_{n,j}\int_{(j-1)/n}^{j/n}f(x){\rm d}x,\quad f\in{\mathcal{F}}_{\delta},$
with $\delta>0$. (Note that we do not assume $f$ to be continuous for $f\in{\mathcal{F}}_{\delta}$. If continuity is assumed, then one can replace $\int_{(j-1)/n}^{j/n}f(x){\rm d}x$ by $f(j/n)/n$ and establish the same limit theorem below.)
The second main result is the following limit theorem.
Theorem 3.1.
Assume $\lim_{n\to\infty}m_{n}/n=\infty$ and $\delta>0$. Then, for all $f\in{\mathcal{F}}_{\delta}$,
$\lim_{n\to\infty}{\mathbb{E}}\exp\left({\rm i}G_{n}(f)\right)={\mathbb{E}}\exp\left({\rm i}G(f)\right)=\exp\left(-\frac{1}{2}{\rm{Cov}}(G(f),G(f))\right).$
In particular, $\{G_{n}(f)\}_{f\in{\mathcal{F}}_{\delta}}\xrightarrow{\textit{f.d.d.}}\{G(f)\}_{f\in{\mathcal{F}}_{\delta}}$ as $n\to\infty$.
Remark 3.2.
We impose the condition in ${\mathcal{F}}_{\delta}$ as it is convenient for establishing our limit theorems. In particular, ${\mathcal{F}}_{\delta}$ includes all smooth functions that decay rapidly at infinity. Lévy’s continuity theorem yields immediately the convergence of $G_{n}$ in the space of tempered distributions. For the generalized process $G$ indexed by $\mathsf{M}$}, Lévy’s continuity theorem applies as long as the space of functions on $\mathsf{M}$} is a nuclear space [12, 34]. We do not pursue to strengthen the convergence in the space of distributions here.
Remark 3.3.
The process $\widehat{S}_{n,j}$ can be thought of the $H=0$ counterpart of the model considered in Enriquez, [11] and Shen et al., [47]. Therein, the density of $q$ is asymptotically equivalent to $Cr^{-2H}$ as $q\downarrow 0$ for some constant $C>0$, and the normalization $1/\sqrt{m_{n}}$ in $\widehat{S}_{n,j}$ is replaced by $1/(n^{H}\sqrt{m_{n}})$, and it was shown that $\{\widehat{S}_{n,\left\lfloor nt\right\rfloor}\}_{t\in[0,1]}\xrightarrow{\textit{f.d.d.}}c\{G^{H}_{t}\}_{t\in[0,1]}$ for some explicit constant $c$. (Therein, there is also no centering in $\widehat{S}_{n,j}$ and so the limit is actually ${\mathbb{B}}^{H}$ instead of $G^{H}$.)
3.1. Proof of Theorem 3.1
Since all $G_{n}(f)$ and $G(f)$ are linear in $f$, it suffices to show
(3.3) $\lim_{n\to\infty}{\mathbb{E}}\exp\left({\rm i}G_{n}(f)\right)={\mathbb{E}}\exp\left({\rm i}G(f)\right)=\exp\left(-\frac{1}{2}{\rm{Cov}}(G(f),G(f))\right),$
for all $f\in{\mathcal{F}}_{\delta},\delta>0$. We start by evaluating ${\mathbb{E}}e^{{\rm i}G_{n}(f)}$. We first introduce some notations. Let $(\eta_{j}^{(r)})_{j\in\mathbb{N}}$ denote i.i.d. Bernoulli random variables with parameter $r\in(0,1)$, and set $\tau_{j}^{(r)}:=\sum_{k=1}^{j}\eta_{k}^{(r)}$. Introduce
$p_{j}(r):=\mathbb{P}\left(\tau_{j}^{(r)}{\rm~is~odd}\right),r\in(0,1).$
Note that here $r$ is a fixed real number. We also let $q$ denote a uniform random variable, and accordingly we write $p_{j}(q)=\mathbb{P}(\tau_{j}^{(q)}{\rm~is~odd}\mid q)$. Write
$\psi_{n,r}(f):=2\sum_{j=1}^{\infty}f_{n,j}\left({{\bf 1}}_{\left\{\tau_{j}^{(r)}{\rm~is~odd}\right\}}-p_{j}(r)\right)\quad\mbox{ with }\quad f_{n,j}:=\int_{(j-1)/n}^{j/n}f(x){\rm d}x.$
Let $\mathcal{X}$ denote a standard Gaussian random variable. Since $G_{n}(f)$ is the sum of $m_{n}$ i.i.d. random variables (recall (3.2)), we have
${\mathbb{E}}e^{{\rm i}G_{n}(f)}$ $=\left({\mathbb{E}}\exp\left(\frac{2{\rm i}}{\sqrt{m_{n}}}\sum_{j=1}^{\infty}f_{n,j}\frac{\mathcal{X}}{q^{1/2}}\left({{\bf 1}}_{\left\{\tau_{j}^{(q)}{\rm~is~odd}\right\}}-p_{j}(q)\right)\right)\right)^{m_{n}}$
$=\left(\int_{0}^{1}{\mathbb{E}}\exp\left(-\frac{1}{2m_{n}r}\psi_{n,r}(f)^{2}\right){\rm d}r\right)^{m_{n}}.$
The goal is to write the above as $(1+v_{n}(f)/m_{n})^{m_{n}}$ with $v_{n}(f)\to-{\rm{Cov}}(G(f),G(f))/2$ as $n\to\infty$. For this purpose, we write
$\int_{0}^{1}{\mathbb{E}}\exp\left(-\frac{1}{2m_{n}r}\psi_{n,r}(f)^{2}\right){\rm d}r=1+\int_{0}^{1}\left({\mathbb{E}}\exp\left(-\frac{1}{2m_{n}r}\psi_{n,r}(f)^{2}\right)-1\right){\rm d}r.$
Fix $\epsilon>0$. In order to prove (3.3) we shall prove
(3.4) $\lim_{\epsilon\downarrow 0}\lim_{n\to\infty}m_{n}\int_{\epsilon/n}^{\epsilon^{-1}/n}\left({\mathbb{E}}\exp\left(-\frac{1}{2m_{n}r}\psi_{n,r}(f)^{2}\right)-1\right){\rm d}r$ $=-\frac{1}{2}{\rm{Cov}}(G(f),G(f)),$
(3.5) $\lim_{\epsilon\downarrow 0}\limsup_{n\to\infty}m_{n}\int_{[0,1]\setminus[\epsilon/n,\epsilon^{-1}/n]}\left({\mathbb{E}}\exp\left(-\frac{1}{2m_{n}r}\psi_{n,r}(f)^{2}\right)-1\right){\rm d}r$ $=0.$
(i) We start by proving (3.4). First,
(3.6) $\int_{\epsilon/n}^{\epsilon^{-1}/n}\left({\mathbb{E}}\exp\left(-\frac{1}{2m_{n}r}\psi_{n,r}(f)^{2}\right)-1\right){\rm d}r\\ =\frac{1}{n}\int_{\epsilon}^{\epsilon^{-1}}\left({\mathbb{E}}\exp\left(-\frac{n}{2m_{n}r}\psi_{n,r/n}(f)^{2}\right)-1\right){\rm d}r\sim-\frac{1}{2m_{n}}\int_{\epsilon}^{\epsilon^{-1}}\frac{1}{r}{\mathbb{E}}\psi_{n,r/n}(f)^{2}{{\rm d}r}.$
In the last step we used the assumption that $m_{n}\gg n$ and that
$\limsup_{n\to\infty}\sup_{r\in[\epsilon,\epsilon^{-1}]}\frac{\psi_{n,r/n}(f)^{2}}{r}<\infty.$
This is because for $f\in{\mathcal{F}}_{\delta}$ we have for all $r>0$,
$|\psi_{n,r}(f)|\leq 2\sum_{j=1}^{\infty}f_{n,j}\leq 2{\left\|f\right\|}_{1}<\infty.$
For each $r>0$ fixed, we shall establish the following.
Lemma 3.4.
For all $r>0,\delta>0$ and $f\in{\mathcal{F}}_{\delta}$,
$\lim_{n\to\infty}{\mathbb{E}}\psi_{n,r/n}(f)^{2}$ $=4{\mathbb{E}}\left(\int_{0}^{\infty}f(s)\left({{\bf 1}}_{\left\{N(rs){\rm~is~odd}\right\}}-\widetilde{p}_{r}(s)\right){\rm d}s\right)^{2}$
$=\int_{0}^{\infty}\int_{0}^{\infty}f(s)f(t)\log\frac{s+t}{|t-s|}{\rm d}s{\rm d}t.$
Proof.
Indeed,
${\mathbb{E}}\psi_{n,r/n}(f)^{2}={\mathbb{E}}\left(2\sum_{j=1}^{\infty}f_{n,j}\left({{\bf 1}}_{\left\{\tau_{j}^{(r/n)}{\rm~is~odd}\right\}}-p_{j}(r/n)\right)\right)^{2}.$
Fix a parameter $K\in\mathbb{N}$, and write
$\psi_{n,r,K}(f):=2\sum_{j=1}^{Kn}f_{n,j}\left({{\bf 1}}_{\left\{\tau_{j}^{(r)}{\rm~is~odd}\right\}}-p_{j}(r)\right),$
and $\widetilde{\psi}_{n,r,K}(f):=\psi_{n,r}(f)-\psi_{n,r,K}(f)$. We first show that
(3.7) $\lim_{n\to\infty}{\mathbb{E}}\psi_{n,r/n,K}(f)^{2}=4{\mathbb{E}}\left(\int_{0}^{K}f(t)\left({{\bf 1}}_{\left\{N_{r}(t){\rm~is~odd}\right\}}-\widetilde{p}_{r}(t)\right){\rm d}t\right)^{2}.$
It suffices to show
(3.8) $\sum_{j=1}^{Kn}f_{n,j}{{\bf 1}}_{\left\{\tau_{j}^{(r/n)}{\rm~is~odd}\right\}}\Rightarrow\int_{0}^{K}f(x){{\bf 1}}_{\left\{N_{r}(x){\rm~is~odd}\right\}}{\rm d}x.$
Then, since the random variables on the left-hand side are all bounded by $K{\left\|f\right\|}_{\infty}$, we have the convergence of all moments, and in particular (3.7) holds.
The convergence (3.8) essentially follows from the well-known Poisson convergence of triangular arrays of i.i.d. Bernoulli random variables
(3.9) $\sum_{j=1}^{Kn}\eta_{j}^{(r/n)}\delta_{j/n}\Rightarrow{\rm PPP}([0,K],r{\rm d}x),$
where the right-hand side denotes the Poisson point process on $[0,K]$ with intensity measure $r{\rm d}x$, and a continuous-mapping argument. Recall that $r>0$ is fixed in (3.8).
First, we re-state the point-process convergence in (3.9) as the convergence of a sequence of random points. Set $\widehat{j}_{n,1}:=\min\{j\in\mathbb{N}:\eta_{j}^{(r/n)}=1\}$,
$\widehat{j}_{n,k+1}:=\min\left\{j>\widehat{j}_{n,k}:\eta_{j}^{(r/n)}=1\right\},k\in\mathbb{N},$
and set
$N_{K,n}:=\max\left\{k\in\mathbb{N}:\widehat{j}_{n,k}\leq Kn\right\}.$
So the point process on the left-hand side of (3.9) is now $\sum_{k=1}^{N_{K,n}}\delta_{\widehat{j}_{n,k}/n}$. Set also $\{\Gamma_{j}\}_{j\in\mathbb{N}}$ be the sequence of consecutive arrival times of the Poisson process $\{N(rt)\}_{t\geq 0}$, and set $N_{K}:=\max\{j\in\mathbb{N}_{0}:\Gamma_{j}\leq K\}$. We then have $N_{K,n}\Rightarrow N_{K}$ and
(3.10) $\left(\widehat{j}_{n,1},\dots,\widehat{j}_{n,N_{K,n}},0,\dots\right)\Rightarrow\left(\Gamma_{1},\dots,\Gamma_{N_{K}},0,\dots\right)$
as weak convergence of random elements in $\{(a_{j})_{j\in\mathbb{N}}:a_{j}=0\mbox{ for $j$ large enough}\}$.
The left-hand side of (3.8) can be re-written as
$F_{K,n}:=\sum_{k=1}^{\lfloor N_{K,n}/2\rfloor}\int^{\widehat{j}_{n,2k}/n}_{\widehat{j}_{n,2k-1}/n}f(t){\rm d}t+{{\bf 1}}_{\left\{N_{K,n}{\rm~is~odd}\right\}}\int^{K}_{\widehat{j}_{n,N_{K,n}}/n}f(t){\rm d}t.$
To show (3.8) now it suffices to show
(3.11) $F_{K,n}\Rightarrow F_{K}:=\sum_{k=1}^{N_{K}/2}\int^{\Gamma_{2k}}_{\Gamma_{2k-1}}f(t){\rm d}t+{{\bf 1}}_{\left\{N_{K}{\rm~is~odd}\right\}}\int_{\Gamma_{N_{K}}}^{K}f(t){\rm d}t.$
The above follows from a continuous mapping theorem applied to (3.10). One way to make this precise is to, by Skorokhod representation theorem, assume that the convergence in (3.10) is further in the almost sure sense, and then examine (3.11) for every $\omega$ such that the sequence convergences. Fix such an $\omega$, and notice that for $n$ large enough $N_{K,n}(\omega)=N_{K}(\omega)$, and then both $F_{K,n}$ and $F_{K}$ are expressed the same continuous function of $N_{K}(\omega)$ random variables. It is now clear that (3.11) converges almost surely. We have thus proved (3.8), and hence (3.7).
To conclude the proof, it remains to show that
$\lim_{K\to\infty}\limsup_{n\to\infty}{\mathbb{E}}\widetilde{\psi}_{n,r/n,K}(f)^{2}=0.$
For this purpose, we could simply bound
${\mathbb{E}}\widetilde{\psi}_{n,r/n,K}(f)^{2}\leq\left(2\sum_{j=Kn+1}^{\infty}f_{n,j}\right)^{2}\leq 4\left(\int_{K}^{\infty}|f(x)|{\rm d}x\right)^{2},$
and the desired result follows from the fact that ${\left\|f\right\|}_{1}<\infty$. ∎
Following Lemma 3.4, we have
$\lim_{n\to\infty}\int_{\epsilon/n}^{\epsilon^{-1}/n}\frac{{\mathbb{E}}\psi_{n,r/n}(f)^{2}}{r}{\rm d}r=\int_{0}^{\infty}\int_{0}^{\infty}f(s)f(t)\int_{\epsilon}^{\epsilon^{-1}}\frac{e^{-2r(t-s)}-e^{-2r(t+s)}}{r}{\rm d}r{\rm d}s{\rm d}t,$
which as $\epsilon\downarrow 0$ converges to ${\rm{Cov}}(G(f),G(f))$ by (2.7). Combining with (3.6), we have thus proved (3.4).
(ii) Now we prove (3.5). applying the simple inequality
$1-{\mathbb{E}}\exp\left(-\frac{1}{2m_{n}r}\psi_{n,r}(f)^{2}\right)\leq\frac{1}{2m_{n}r}{\mathbb{E}}\psi_{n,r}(f)^{2},$
it suffices to show
(3.12) $\lim_{\epsilon\downarrow 0}\limsup_{n\to\infty}$ $\int_{0}^{\epsilon}\frac{1}{r}{\mathbb{E}}\psi_{n,r/n}(f)^{2}{\rm d}r=0,$
(3.13) $\lim_{\epsilon\downarrow 0}\limsup_{n\to\infty}$ $\int_{\epsilon^{-1}}^{n}\frac{1}{r}{\mathbb{E}}\psi_{n,r/n}(f)^{2}{\rm d}r=0.$
We first show (3.13). Note that, for $j\leq j^{\prime}$,
$\mathbb{P}$ $\left(\tau_{j}^{(r)}{\rm~is~odd},\tau_{j^{\prime}}^{(r)}{\rm~is~odd}\right)-\mathbb{P}\left(\tau_{j}^{(r)}{\rm~is~odd}\right)\mathbb{P}\left(\tau_{j^{\prime}}^{(r)}{\rm~is~odd}\right)$
$=\frac{1}{4}\left(\left(1-(1-2r)^{j}\right)\left(1+(1-2r)^{j^{\prime}-j}\right)-\left(1-(1-2r)^{j}\right)\left(1-(1-2r)^{j^{\prime}}\right)\right)$
$=\frac{1}{4}\left(1-(1-2r)^{j}\right)\left((1-2r)^{j^{\prime}-j}+(1-2r)^{j^{\prime}}\right)=\frac{1}{4}\left(1-(1-2r)^{2j}\right)(1-2r)^{j^{\prime}-j}.$
Therefore, we have
${\mathbb{E}}\psi_{n,r/n}(f)^{2}$ $\leq 8\sum_{j=1}^{\infty}\sum_{j^{\prime}=j}^{\infty}\left|f_{n,j}f_{n,j^{\prime}}\right|\frac{1}{4}\left(1-\left(1-\frac{2r}{n}\right)^{2j}\right)\left(1-\frac{2r}{n}\right)^{j^{\prime}-j}$
(3.14) $=8\sum_{j=1}^{\infty}\sum_{k=0}^{\infty}\left|f_{n,j}f_{n,j+k}\right|\frac{1}{4}\left(1-\left(1-\frac{2r}{n}\right)^{2j}\right)\left(1-\frac{2r}{n}\right)^{k}$
$\leq 2\sum_{j=1}^{\infty}|f_{n,j}|\sum_{k=0}^{\infty}\frac{{\left\|f\right\|}_{\infty}}{n}\left(1-\frac{2r}{n}\right)^{k}=\frac{{\left\|f\right\|}_{\infty}{\left\|f\right\|}_{1}}{r}.$
Thus,
$\int_{\epsilon^{-1}}^{n}\frac{{\mathbb{E}}\psi_{n,r/n}(f)^{2}}{r}{\rm d}r\leq C\int_{\epsilon^{-1}}^{\infty}\frac{1}{r^{2}}{\rm d}r=C\epsilon.$
The desired (3.13) now follows.
We next show (3.12). Break the expression (3.14) into two parts depending on $j/n\leq 1/r$ and $j/n>1/r$. First,
$8\sum_{j=1}^{\left\lfloor n/r\right\rfloor}\left|f_{n,j}\right|\sum_{k=0}^{\infty}\left|f_{n,j+k}\right|\frac{1}{4}\left(1-\left(1-\frac{2r}{n}\right)^{2j}\right)\left(1-\frac{2r}{n}\right)^{k}$
$\leq 8r\sum_{j=1}^{\left\lfloor n/r\right\rfloor}\frac{j}{n}\left|f_{n,j}\right|{\left\|f\right\|}_{1}\leq 8r^{\delta}\sum_{j=1}^{\left\lfloor n/r\right\rfloor}\left(\frac{j}{n}\right)^{\delta}\left|f_{n,j}\right|{\left\|f\right\|}_{1}\leq Cr^{\delta}\int_{0}^{\infty}s^{\delta}|f(s)|{\rm d}s{\left\|f\right\|}_{1},$
where in the first step we used $1-x^{p}=(1-x)(1+x+\cdots+x^{p-1})\leq p(1-x)$ for $x\in[-1,1],p\in\mathbb{N}$. Also,
$8\sum_{j=\left\lfloor n/r\right\rfloor+1}^{\infty}\left|f_{n,j}\right|\sum_{k=0}^{\infty}\left|f_{n,j+k}\right|\frac{1}{4}\left(1-\left(1-\frac{2r}{n}\right)^{2j}\right)\left(1-\frac{2r}{n}\right)^{k}$
$\leq 8\sum_{j=\left\lfloor n/r\right\rfloor+1}^{\infty}\left|f_{n,j}\right|{\left\|f\right\|}_{1}\leq 8r^{\delta}\sum_{j=\left\lfloor n/r\right\rfloor+1}^{\infty}\left(\frac{j}{n}\right)^{\delta}\left|f_{n,j}\right|{\left\|f\right\|}_{1}\leq Cr^{\delta}\int_{0}^{\infty}s^{\delta}|f(s)|{\rm d}s{\left\|f\right\|}_{1}.$
That is, we have shown that ${\mathbb{E}}\psi_{n,r/n}(f)^{2}\leq Cr^{\delta}$, whence
$\int_{0}^{\epsilon}{\mathbb{E}}\psi_{n,r/n}(f)^{2}\frac{{\rm d}r}{r}\leq C\epsilon^{\delta},$
and hence (3.13) follows. This completes the proof of Theorem 3.1.
4. Log-correlated bi-fractional Brownian motions indexed by metric spaces
We introduce a family of log-correlated generalized Gaussian processes indexed by metric spaces $({\mathsf{M}},{\mathsf{d}})$ under the Assumption 1.2 and their stable extensions in a unified framework. We also provide a central limit theorem.
4.1. A stochastic-integral representation
Let $(E,{\mathcal{E}},\mu)$ and $\{A_{x}\}_{x\in{\mathsf{M}}}\subset{\mathcal{E}}$ be as in Assumption 1.2. Let ${\mathbb{M}}_{p}(E)$ denote the space of Radon point measures on $E$, and $\mathcal{M}_{p}(E)$ its Borel $\sigma$-algebra (see e.g. [40]). For notational convenience, we write
$\beta=2H\in(0,1]$
in this section. For each $r>0$, let $\mathbb{P}_{\beta,r}^{\prime}\equiv\mathbb{P}_{\beta,r,\mu}^{\prime}$ denote the probability measure on $({\mathbb{M}}_{p}(E),\mathcal{M}_{p}(E))$ induced by the Poisson point process on $(E,{\mathcal{E}})$ with random intensity measure $\mathcal{S}_{\beta,r}\cdot\mu$, where $\mathcal{S}_{\beta,r}$ is a totally skewed $\beta$-stable random variable with
${\mathbb{E}}\exp(-\theta\mathcal{S}_{\beta,r})=\exp\left(-r\theta^{\beta}\right),\theta>0.$
(That is, $\mathcal{S}_{\beta,r}$ is a $\beta$-stable subordinator evaluated at time $r$.) Note that the special case $\mathcal{S}_{1,r}=r$ (and this section might be easier to read first assuming $\beta=1$). In particular, letting $m$ denote an element in ${\mathbb{M}}_{p}(E)$,
$\mathbb{P}^{\prime}_{\beta,r}(m(A)=k)$ $\equiv\mathbb{P}^{\prime}_{\beta,r}\left(\left\{m\in\mathbb{M}_{p}(E):m(A)=k\right\}\right)$
$={\mathbb{E}}\left(\frac{(\mathcal{S}_{\beta,r}\mu(A))^{k}}{k!}e^{-\mathcal{S}_{\beta,r}\mu(A)}\right),k\in\mathbb{N}_{0},$
for all $A\in{\mathcal{E}},\mu(A)<\infty$. Set
$\widetilde{p}_{\beta,r}(x):=\mathbb{P}^{\prime}_{\beta,r}\left(m(A){\rm~is~odd}\right)={\mathbb{E}}\left(\frac{1}{2}\left(1-e^{-2\mathcal{S}_{\beta,r}\mu(A_{x})}\right)\right)=\frac{1}{2}\left(1-e^{-(2\mu(A_{x}))^{\beta}r}\right),x\in{\mathsf{M}}.$
Let $\lambda$ be a Radon measure on $({\mathsf{M}},{\mathsf{d}})$. Introduce, for $\alpha\in(0,2]$,
$G_{\alpha,\beta}(f):=\int_{\mathbb{R}_{+}\times{\mathbb{M}}_{p}(E)}\int_{x\in{\mathsf{M}}}\left({{\bf 1}}_{\left\{m(A_{x}){\rm~is~odd}\right\}}-\widetilde{p}_{\beta,r}(x)\right)f(x)\lambda({\rm d}x)M_{\alpha,\beta}({\rm d}r{\rm d}m),$
where $M_{\alpha,\beta}$ is a symmetric $\alpha$-stable random measure on $\mathbb{R}_{+}\times{\mathbb{M}}_{p}(E)$ with control measure $2r^{-1}{\rm d}r{\rm d}\mathbb{P}_{\beta,r}^{\prime})$, and $f\in{\mathcal{F}}_{\delta,\delta^{\prime},\beta}({\mathsf{M}})$ defined as follows.
Assumption 4.1.
With $\delta,\delta^{\prime}>0,\beta\in(0,1]$, we let ${\mathcal{F}}_{\delta,\delta^{\prime},\beta}({\mathsf{M}})$ denote the class of all bounded and measurable functions $f:{\mathsf{M}}\to\mathbb{R}$ (measurable with respect to the Borel $\sigma$-algebra induced by $\mathsf{d}$}) satisfying the following:
  1. (i)
    ${\left\|f\right\|}_{1}:=\int_{\mathsf{M}}|f(x)|\lambda({\rm d}x)<\infty$,
  2. (ii)
    ${\left\|f\right\|}_{1,\delta}:=\int_{\mathsf{M}}\mu(A_{x})^{\delta}|f(x)|\lambda({\rm d}x)<\infty$,
  3. (iii)
    $\int_{\mathsf{M}}\int_{\mathsf{M}}e^{-{\mathsf{d}}^{\beta}(x,y)r}|f(x)f(y)|\lambda({\rm d}x)\lambda({\rm d}y)\leq Cr^{-\delta^{\prime}}$ for all $r>0$.
It is useful to recall that when $\mu(A_{o})=0$ for some $o\in{\mathsf{M}}$ (usually playing the role of the origin), we have $\mu(A_{x})={\mathsf{d}}(o,x)$. In particular, the above assumption can be viewed as a generalization of the assumption on $f$ in (2.12).
The following is the main result for log-correlated bi-fractional Brownian motions and their stable extensions indexed by metric spaces, where $\Gamma^{(\beta)}(x,y)$ is the same as in Theorem 1.3.
Theorem 4.2.
For each $\alpha\in(0,2],\delta,\delta^{\prime}>0$, $\{G_{\alpha,\beta}(f)\}_{f\in{\mathcal{F}}_{\delta,\delta^{\prime},\beta}({\mathsf{M}})}$ is a family of symmetric $\alpha$-stable random variables with
(4.1) ${\mathbb{E}}\exp\left({\rm i}G_{\alpha,\beta}(f)\right)=\exp\left(-2\int_{0}^{\infty}{\mathbb{E}}^{\prime}_{\beta,r}\left|\int_{\mathsf{M}}\left({{\bf 1}}_{\left\{m(A_{x}){\rm~is~odd}\right\}}-\widetilde{p}_{\beta,r}(x)\right)f(x)\lambda({\rm d}x)\right|^{\alpha}\frac{{\rm d}r}{r}\right).$
In particular, when $\alpha=2$, this is a family of centered Gaussian random variables with
(4.2) ${\rm{Cov}}(G_{2,\beta}(f),G_{2,\beta}(g))=\int_{\mathsf{M}}\int_{\mathsf{M}}f(x)g(y)\Gamma^{(\beta)}(x,y)\lambda({\rm d}x)\lambda({\rm d}y),$
for all $f,g\in{\mathcal{F}}_{\delta,\delta^{\prime},\beta}({\mathsf{M}})$, where
(4.3) $\Gamma^{(\beta)}(x,y)=\log\frac{\mu^{\beta}(A_{x})+\mu^{\beta}(A_{y})}{{\mathsf{d}}^{\beta}(x,y)}=4\int_{0}^{\infty}\frac{\Gamma_{r}^{(\beta)}(x,y)}{r}{\rm d}r,$
with
(4.4) $\Gamma_{r}^{(\beta)}(x,y)\\ :={\mathbb{E}}^{\prime}_{\beta,r}\left(\left({{\bf 1}}_{\left\{m(A_{x}){\rm~is~odd}\right\}}-\mathbb{P}_{\beta,r}^{\prime}\left(m(A_{x}){\rm~is~odd})\right)\right)\left({{\bf 1}}_{\left\{m(A_{y}){\rm~is~odd}\right\}}-\mathbb{P}_{\beta,r}^{\prime}\left(m(A_{y}){\rm~is~odd}\right)\right)\right).$
Proof.
This time, the identity (4.1) again follows from the property of stochastic integrals, and we only need to show the finiteness of the right-hand side of (4.1). Write
(4.5) $\Psi_{\alpha,\beta,f}(r)={\mathbb{E}}^{\prime}_{\beta,r}\left|\int_{\mathsf{M}}\left({{\bf 1}}_{\left\{m(A_{x}){\rm~is~odd}\right\}}-\widetilde{p}_{r}(x)\right)f(x)\lambda({\rm d}x)\right|^{\alpha}.$
The goal is to show
(4.6) $\int_{0}^{\infty}\frac{\Psi_{\alpha,\beta,f}(r)}{r}{\rm d}r<\infty.$
We first deal with $\alpha=2$. We compute
$\mathbb{P}^{\prime}_{\beta,r}$ $\left(m(A_{x}){\rm~is~odd},m(A_{y}){\rm~is~odd}\right)$
$=\mathbb{P}^{\prime}_{\beta,r}\left(m(A_{x}\setminus A_{y}){\rm~is~odd}\right)\mathbb{P}^{\prime}_{\beta,r}\left(m(A_{y}\setminus A_{x}){\rm~is~odd}\right)\mathbb{P}^{\prime}_{\beta,r}\left(m(A_{x}\cap A_{y}){\rm~is~even}\right)$
$\quad+\mathbb{P}^{\prime}_{\beta,r}\left(m(A_{x}\setminus A_{y}){\rm~is~even}\right)\mathbb{P}^{\prime}_{\beta,r}\left(m(A_{y}\setminus A_{x}){\rm~is~even}\right)\mathbb{P}^{\prime}_{\beta,r}\left(m(A_{x}\cap A_{y}){\rm~is~odd}\right)$
$={\mathbb{E}}^{\prime}_{\beta,r}\left(\frac{1}{4}\left(1-e^{-2\mathcal{S}_{\beta,r}\mu(A_{x})}-e^{-2\mathcal{S}_{\beta,r}\mu(A_{y})}+e^{-2\mathcal{S}_{\beta,r}\mu(A_{x}\Delta A_{y})}\right)\right)$
$=\frac{1}{4}\left(1-e^{-(2\mu(A_{x}))^{\beta}r}-e^{-(2\mu(A_{y}))^{\beta}r}+e^{-(2\mu(A_{x}\Delta A_{y}))^{\beta}r}\right).$
Then, taking $\Gamma^{(\beta)}_{r}(x,y)$ as given in (4.4), we have
(4.7) $\Gamma_{r}^{(\beta)}(x,y)$ $=\mathbb{P}^{\prime}_{\beta,r}\left(m(A_{x}){\rm~is~odd},m(A_{y}){\rm~is~odd}\right)-\mathbb{P}^{\prime}_{\beta,r}\left(m(A_{x}){\rm~is~odd}\right)\mathbb{P}^{\prime}_{\beta,r}\left(m(A_{y}){\rm~is~odd}\right)$
$=\frac{1}{4}\left(e^{-(2\mu(A_{x}\Delta A_{y}))^{\beta}r}-e^{-((2\mu(A_{x}))^{\beta}+(2\mu(A_{y}))^{\beta})r}\right)\geq 0.$
Therefore, we have
$\Psi_{2,\beta,f}(r)$
$={\mathbb{E}}^{\prime}_{\beta,r}\int_{\mathsf{M}}\int_{\mathsf{M}}\left({{\bf 1}}_{\left\{m(A_{x}){\rm~is~odd}\right\}}-\widetilde{p}_{\beta,r}(A_{x})\right)\left({{\bf 1}}_{\left\{m(A_{y}){\rm~is~odd}\right\}}-\widetilde{p}_{\beta,r}(A_{y})\right)f(x)f(y)\lambda({\rm d}x)\lambda({\rm d}y)$
$=\int_{\mathsf{M}}\int_{\mathsf{M}}\Gamma_{r}^{(\beta)}(x,y)f(x)f(y)\lambda({\rm d}x)\lambda({\rm d}y)$
$=\frac{1}{4}\int_{\mathsf{M}}\int_{\mathsf{M}}\left(e^{-(2\mu(A_{x}\Delta A_{y}))^{\beta}r}-e^{-((2\mu(A_{x}))^{\beta}+(2\mu(A_{y}))^{\beta})r}\right)f(x)f(y)\lambda({\rm d}x)\lambda({\rm d}y).$
The first equality is the definition, the second follows from Fubini’s theorem (since ${\left\|f\right\|}_{1}<\infty$), and the third follows from (4.7).
Next, we prove
(4.8) $\Psi_{2,\beta,f}(r)\leq C(r^{\delta}\wedge r^{-\delta^{\prime}})\mbox{ for all }r>0,$
for all $f\in{\mathcal{F}}_{\delta,\delta^{\prime},\beta}({\mathsf{M}})$ (the constant $C$ depends on $f$). For the sake of simplicity, assume that $f\geq 0$. We break the outer integral into two parts and proceed as follows. First,
$\int_{\mu(A_{x})>1/r}$ $\int_{\mathsf{M}}f(x)f(y)\left(e^{-(2\mu(A_{x}\Delta A_{y}))^{\beta}r}-e^{-((2\mu(A_{x}))^{\beta}+(2\mu(A_{y}))^{\beta})r}\right)\lambda({\rm d}x)\lambda({\rm d}y)$
$\leq\int_{\mu(A_{x})>1/r}\int_{\mathsf{M}}f(x)f(y)\lambda({\rm d}x)\lambda({\rm d}y)=\int_{\mu(A_{x})>1/r}f(x)\lambda({\rm d}x){\left\|f\right\|}_{1}$
$\leq r^{\delta}{\left\|f\right\|}_{1}\int_{\mu(A_{x})>1/r}\mu(A_{x})^{\delta}f(x)\lambda({\rm d}x)\leq C{\left\|f\right\|}_{1}{\left\|f\right\|}_{1,\delta}r^{\delta}.$
Second, using
$e^{-(2\mu(A_{x}\Delta A_{y}))^{\beta}r}-e^{-((2\mu(A_{x}))^{\beta}+(2\mu(A_{y}))^{\beta})r}$ $\leq 1-e^{-2^{\beta}(\mu^{\beta}(A_{x})+\mu^{\beta}(A_{y})-\mu^{\beta}(A_{x}\Delta A_{y}))r}$
$\leq 2^{\beta}(\mu^{\beta}(A_{x})+\mu^{\beta}(A_{y})-\mu^{\beta}(A_{x}\Delta A_{y}))r$
$\leq 2^{\beta+1}\mu^{\beta}(A_{x})r,$
we have
$\int_{\mu(A_{x})\leq 1/r}$ $\int_{\mathsf{M}}f(x)f(y)\left(e^{-(2\mu(A_{x}\Delta A_{y}))^{\beta}r}-e^{-((2\mu(A_{x}))^{\beta}+(2\mu(A_{y}))^{\beta})r}\right)\lambda({\rm d}x)\lambda({\rm d}y)$
$\leq{\left\|f\right\|}_{1}\int_{\mu(A_{x})\leq 1/r}2^{\beta+1}r\mu^{\beta}(A_{x})f(x)\lambda({\rm d}x)$
$\leq Cr\int_{\mu(A_{x})\leq 1/r}(1/r)^{\beta(1-\widetilde{\delta})}\mu^{\beta\widetilde{\delta}}(A_{x})f(x)\lambda({\rm d}x)\leq Cr^{1-\beta+\beta\widetilde{\delta}}{\left\|f\right\|}_{1,\beta\widetilde{\delta}}$,
for $\widetilde{\delta}\in(0,1)$. Taking $\widetilde{\delta}=\delta/\beta$, the above is bounded by $Cr^{1-\beta+\delta}\leq Cr^{\delta}$ for $r\in[0,1]$. We have thus proved $\Psi_{2,\beta,f}(r)\leq Cr^{\delta}$ for $r\in[0,1]$. Alternatively, using $e^{-(2\mu(A_{x}\Delta A_{y}))^{\beta}r}-e^{-((2\mu(A_{x}))^{\beta}+(2\mu(A_{y}))^{\beta})r}\leq e^{-2r\mu(A_{x}\Delta A_{y})}=e^{-2^{\beta}{\mathsf{d}}^{\beta}(x,y)r}$, we have
$|\Psi_{2,\beta,f}(r)|\leq\int_{\mathsf{M}}\int_{\mathsf{M}}e^{-2^{\beta}{\mathsf{d}}^{\beta}(x,y)r}f(x)f(y)\lambda({\rm d}y)\lambda({\rm d}x)\leq Cr^{-\delta^{\prime}},$
where the last step follows from (iii) of Assumption 4.1. We have thus proved (4.8).
Now (4.6) follows with $\alpha=2$. For $\alpha\in(0,2)$, using $\Psi_{\alpha,\beta,f}(r)\leq\Psi_{2,\beta,f}(r)^{\alpha/2}$, (4.6) remains to hold. We have thus proved that (4.1) holds and the right-hand side is finite.
In particular for the case $\alpha=2$, (4.1), (4.7) and Gamma integral yield
${\rm{Cov}}(G_{2,\beta}(f),G_{2,\beta}(f))$ $=4\int_{0}^{\infty}\int_{\mathsf{M}}\int_{\mathsf{M}}f(x)f(y)\Gamma_{r}^{(\beta)}(x,y)\lambda({\rm d}x)\lambda({\rm d}y)\frac{{\rm d}r}{r}$
$=\int_{\mathsf{M}}\int_{\mathsf{M}}f(x)f(y)\log\frac{\mu^{\beta}(A_{x})+\mu^{\beta}(A_{y})}{\mu^{\beta}(A_{x}\Delta A_{y})}\lambda({\rm d}x)\lambda({\rm d}y),$
where we used Frullani integral again. By Assumption 1.2, $\mu(A_{x}\Delta A_{y})={\mathsf{d}}(x,y)$. The stated formula (4.2) for the covariance function now follows. ∎
4.2. A central limit theorem
We consider another aggregated model combining ideas from [11] and [14] that scales to the generalized process just introduced. We only illustrate the case $H=1/2$.
Fix $\alpha\in(0,2]$. Let $\mathcal{X}^{(\alpha)}$ be a symmetric $\alpha$-stable random variable determined by
${\mathbb{E}}\exp\left({\rm i}\theta\mathcal{X}^{(\alpha)}\right)=\exp\left(-|\theta|^{\alpha}\right).$
Let $q$ be a uniform random variable over $(0,1)$ and given $q$ and $n\in\mathbb{N}$, let $N_{n}^{(q)}$ denote a Poisson point process on $(E,{\mathcal{E}})$ with intensity measure $nq\cdot\mu$. Assume $\mathcal{X}^{(\alpha)}$ is independent from $q$ and $N_{n}^{(q)}$. Let $(\mathcal{X}_{i}^{(\alpha)},q_{i},N_{n,i}^{(i)})_{i\in\mathbb{N}}$ be i.i.d. copies of $(\mathcal{X}^{(\alpha)},q,N_{n}^{(q)})$. Set
$G_{\alpha,n}(f):=\frac{1}{m_{n}^{1/\alpha}}\sum_{i=1}^{m_{n}}\frac{\mathcal{X}_{i}^{(\alpha)}}{q_{i}^{1/\alpha}}\int_{\mathsf{M}}\left({{\bf 1}}_{\left\{N_{n,i}^{(q_{i})}(A_{x}){\rm~is~odd}\right\}}-\mathbb{P}\left(N_{n,i}^{(q_{i})}(A_{x}){\rm~is~odd}\;\middle|\;q_{i}\right)\right)f(x)\lambda({\rm d}x),$
for all $f\in{\mathcal{F}}_{\delta,\delta^{\prime},1}({\mathsf{M}})$. This represents an aggregated model based on the triangular arrays of i.i.d. point processes $(N_{n,i}^{(q_{i})})_{n\in\mathbb{N},i\in\mathbb{N}}$.
Theorem 4.3.
Under the assumption $m_{n}\gg n$, we have
$\lim_{n\to\infty}{\mathbb{E}}\exp\left({\rm i}G_{\alpha,n}(f)\right)={\mathbb{E}}\exp\left({\rm i}G_{\alpha}(f)\right),\mbox{ for all }f\in{\mathcal{F}}_{\delta,\delta^{\prime},1}({\mathsf{M}}).$
Proof.
Introduce $N_{n}^{(r)}$ with $r\in(0,1)$ as a fixed number similarly as $N_{n}^{(q)}$. Write $\widetilde{p}_{n,r}(A)=(1/2)(1-e^{-2nr\mu(A)})$. We begin by writing
${\mathbb{E}}e^{{\rm i}G_{\alpha,n}(f)}$ $=\left({\mathbb{E}}\exp\left(\frac{{\rm i}}{m_{n}^{1/\alpha}}\frac{\mathcal{X}^{(\alpha)}}{q^{1/\alpha}}\int_{\mathsf{M}}\left({{\bf 1}}_{\left\{N_{n}^{(q)}(A_{x}){\rm~is~odd}\right\}}-\widetilde{p}_{n,q}(A_{x})\right)f(x)\lambda({\rm d}x)\right)\right)^{m_{n}}$
$=\left(\int_{0}^{1}{\mathbb{E}}\exp\left(-\frac{1}{m_{n}r}\left|\int_{\mathsf{M}}\left({{\bf 1}}_{\left\{N_{n}^{(r)}(A_{x}){\rm~is~odd}\right\}}-\widetilde{p}_{n,r}(A_{x})\right)f(x)\lambda({\rm d}x)\right|^{\alpha}\right){\rm d}r\right)^{m_{n}}$
$=\left(\int_{0}^{n}{\mathbb{E}}\exp\left(-\frac{n}{m_{n}r}\left|\int_{\mathsf{M}}\left({{\bf 1}}_{\left\{N_{1}^{(r)}(A_{x}){\rm~is~odd}\right\}}-\widetilde{p}_{1,r}(A_{x})\right)f(x)\lambda({\rm d}x)\right|^{\alpha}\right)\frac{{\rm d}r}{n}\right)^{m_{n}},$
where in the last step we first applied change of variables $r\mapsto nr$ and then used the fact that $N_{n}^{(r/n)}\stackrel{{\scriptstyle d}}{{=}}N_{1}^{(r)}$ and $\widetilde{p}_{n,r/n}(x)=\widetilde{p}_{1,r}(x)=\widetilde{p}_{r}(x)$. That is, with
$\widehat{\Psi}_{\alpha,f}(r)=\left|\int_{\mathsf{M}}\left({{\bf 1}}_{\left\{N_{1}^{(r)}(A_{x}){\rm~is~odd}\right\}}-\widetilde{p}_{1,r}(A_{x})\right)f(x)\lambda({\rm d}x)\right|^{\alpha},$
we have shown
${\mathbb{E}}\exp\left({\rm i}G_{\alpha,n}(f)\right)=\left(1-\int_{0}^{n}{1-{\mathbb{E}}\exp\left(-\frac{n}{m_{n}r}\widehat{\Psi}_{\alpha,f}(r)\right)}\frac{{\rm d}r}{n}\right)^{m_{n}}.$
It suffices to show
(4.9) $\int_{0}^{n}{1-{\mathbb{E}}\exp\left(-\frac{n}{m_{n}r}\widehat{\Psi}_{\alpha,f}(r)\right)}\frac{{\rm d}r}{n}\sim\frac{1}{m_{n}}\int_{0}^{\infty}\frac{{\mathbb{E}}\widehat{\Psi}_{\alpha,f}(r)}{r}{\rm d}r$
as $n\to\infty$.
Fix $\epsilon\in(0,1)$. First we have
(4.10) $\int_{\epsilon}^{\epsilon^{-1}}{1-{\mathbb{E}}\exp\left(-\frac{n}{m_{n}r}\widehat{\Psi}_{\alpha,f}(r)\right)}\frac{{\rm d}r}{n}\sim\frac{1}{m_{n}}\int_{\epsilon}^{\epsilon^{-1}}\frac{{\mathbb{E}}\widehat{\Psi}_{\alpha,f}(r)}{r}{\rm d}r.$
This step follows from the bounded convergence theorem with the observation $|\widehat{\Psi}_{\alpha,f}(r)|\leq(\int_{\mathsf{M}}|f(x)|\lambda({\rm d}x))^{\alpha}<\infty$ almost surely and the assumption $m_{n}\gg n$. Second, we have
(4.11) $\int_{[0,n]\setminus[\epsilon,\epsilon^{-1}]}{1-{\mathbb{E}}\exp\left(-\frac{n}{m_{n}r}\widehat{\Psi}_{\alpha,f}(r)\right)}\frac{{\rm d}r}{n}\leq\frac{1}{m_{n}}\int_{[0,\epsilon]\cup[\epsilon^{-1},\infty)}\frac{{\mathbb{E}}\widehat{\Psi}_{\alpha,f}(r)}{r}{\rm d}r,$
and (recall $\Psi_{\alpha,1,f}(r)$ in (4.5))
(4.12) ${\mathbb{E}}\widehat{\Psi}_{\alpha,f}(r)=\Psi_{\alpha,1,f}(r)\leq\left(\Psi_{2,1,f}(r)\right)^{\alpha/2}\leq C(r^{\delta}\wedge r^{-\delta^{\prime}})^{\alpha/2},$
for some constant $C$ depending on $f\in{\mathcal{F}}_{\delta,\delta^{\prime},1}({\mathsf{M}})$, where the last step was shown in (4.8). Now, combining (4.10), (4.11) and (4.12) and letting $\epsilon\downarrow 0$, the above estimates yield (4.9). ∎
Appendix A An example by Iksanov and Kotelnikova
Iksanov and Kotelnikova, [22] considered a limit theorem for the Karlin model in the regime that the sampling frequencies $(p_{j})_{j\in\mathbb{N}}$ decay faster than any polynomial rate (recall that for the Karlin model with $\alpha\in(0,1)$, $p_{j}\sim Cj^{-1/\alpha}$). Namely, they considered the case that for $\rho(t):=|\{k\in\mathbb{N}:p_{k}\geq 1/t\}|$, that there exists a slowly varying function $g$ at infinity with $g(x)\to\infty$ as $x\to\infty$ such that $\lim_{x\to\infty}(\rho(\lambda t)-\rho(t))/g(t)=\log\lambda$ for all $\lambda>0$. We omit the description of the model (they actually investigated a sophisticated hierarchical Karlin model and the discussions here concern only their results at the first level), but focus on the limit processes.
In particular, they investigated the scaling limit of the $j$-occupancy process. That is, the counting process for the number of urns containing exactly $j$ after $n$ rounds. The limit process [22, Eq. (4)] is denoted by $(X_{j}(u))_{u\in\mathbb{R}}$ below. The parameterization is a little unconventional as the index set is $\mathbb{R}$, although we shall eventually see that applying a time-change $u\mapsto t=\log u\in(0,\infty)$ we can relate $(X_{j})_{j\in\mathbb{N}}$ to the process of interest here $G^{1/2}$.
Stochastic-integral representations for their limit processes were provided in [22]. Some preparations are needed for the Gaussian random measures involved. Let $\mathbb{P}_{U^{\infty}}$ denote the uniform probability measure on $0,1]^{\mathbb{N}}$ (the distribution of i.i.d. uniform random variables). Let $W_{\infty}$ denote a Gaussian random measure on $\mathbb{R}\times[0,1]^{\mathbb{N}}$ with intensity measure ${\rm d}x{\rm d}\mathbb{P}_{U^{\infty}}$. For each $j\in\mathbb{N}$, let $W_{j}$ denote $W_{\infty}$ restricted to $\mathbb{R}\times[0,1]^{j}$, that is,
$W_{j}({\rm d}x,{\rm d}y_{1},\dots,{\rm d}y_{j}):=W_{\infty}\left({\rm d}x,{\rm d}y_{1},\dots,{\rm d}y_{j},[0,1]^{\mathbb{N}}\right),x\in\mathbb{R},y_{i}\in[0,1],i=1,\dots,j.$
Then, the limit $j$-occupancy processes have the following stochastic-integral representation
(A.1) $\left(X_{j}(u)\right)_{u\in\mathbb{R},j\in\mathbb{N}}\\ =\left(\int_{\mathbb{R}\times[0,1]^{j+1}}\left({{\bf 1}}_{\left\{y_{1}\cdots y_{j+1}<\psi_{0}(e^{-(x-u)})<y_{1}\cdots y_{j}\right\}}-\psi_{j}(e^{-(x-u)})\right)W_{j+1}({\rm d}x{\rm d}y_{1}\cdots{\rm d}y_{j+1})\right)_{u\in\mathbb{R},j\in\mathbb{N}},$
where $\psi_{j}(x)=\mathbb{P}(N_{x}=j)$ and $N_{x}$ is a Poisson random variable with parameter $x>0$ and $j\in\mathbb{N}_{0}$.
The processes $(X_{j})_{j\in\mathbb{N}}$ also have the following representation. Let $(\Omega^{\prime},{\mathcal{F}}^{\prime},\mathbb{P}^{\prime})$ denote a probability space and $M$ a Gaussian random measure on $(\Omega^{\prime},{\mathcal{F}}^{\prime})$ with control measure $r^{-1}{\rm d}r{\rm d}\mathbb{P}^{\prime}$, and $N^{\prime}$ is a standard Poisson process on $(\Omega^{\prime},\mathbb{P}^{\prime})$.
Lemma A.1.
We have the following equivalent representation of (A.1):
$\left(X_{j}(\log t)\right)_{t\in(0,\infty),j\in\mathbb{N}}\stackrel{{\scriptstyle d}}{{=}}\left(\int_{\mathbb{R}\times\Omega^{\prime}}\left({{\bf 1}}_{\left\{N^{\prime}(rt)=j\right\}}-\mathbb{P}(N^{\prime}(rt)=j)\right)M({\rm d}r{\rm d}\omega^{\prime})\right)_{t\in(0,\infty),j\in\mathbb{N}}.$
Proof.
We first re-write the right-hand side of (A.1) in a more probabilistic way. We can view $\widetilde{\Omega}=[0,1]^{\mathbb{N}}$ as a probability space and then $Y_{j}(\widetilde{\omega})\equiv Y_{j}((y_{k})_{k\in\mathbb{N}}):=y_{j}$ as i.i.d. uniform random variables with respect to the law $\mathbb{P}_{U^{\infty}}$. Note also that then $E_{j}:=-\log Y_{j}$ are i.i.d. standard exponential random variables. Then, the event in the indicator function in the integrand of the right-hand side of (A.1) becomes
$\left\{\prod_{k=1}^{j+1}Y_{k}<\exp\left(-e^{-(x-u)}\right)<\prod_{k=1}^{j}Y_{k}\right\}$ $=\left\{\sum_{i=1}^{j+1}(-\log Y_{i})>e^{-(x-u)}>\sum_{i=1}^{j}(-\log Y_{i})\right\}$
$=\left\{\sum_{i=1}^{j}E_{i}<e^{-(x-u)}<\sum_{i=1}^{j+1}E_{i}\right\}=\left\{N(e^{-(x-u)})=j\right\},$
where we also set $N(t)=\max\{j\in\mathbb{N}:\sum_{i=1}^{j}E_{i}\leq t\}$ and $N(t)=0$ if the maximum is over an empty set so that $(N(t))_{t\geq 0}$ with respect to $\mathbb{P}_{U^{\infty}}$ is a standard Poisson process. That is, we have shown that (A.1) can be re-written as
$\left(X_{j}(u)\right)_{u\in\mathbb{R},j\in\mathbb{N}}\\ =\left(\int_{\mathbb{R}\times[0,1]^{\mathbb{N}}}\left({{\bf 1}}_{\left\{N(e^{-(x-u)})=j\right\}}-\mathbb{P}_{U_{\infty}}\left(N\left(e^{-(x-u)}\right)=j\right)\right)W_{\infty}({\rm d}x{\rm d}{\boldsymbol{y}})\right)_{u\in\mathbb{R},j\in\mathbb{N}}.$
Applying the change of variables $r=e^{-x}$ (so ${\rm d}x=r^{-1}{\rm d}r$ and the corresponding domain becomes $r\in(0,\infty)$ and also $t=e^{u}$, we have proved the stated result. ∎
Then, with $\widetilde{X}_{j}(t)=X_{j}(\log t),t\in(0,\infty)$, $\sum_{\ell=1}^{\infty}\widetilde{X}_{2j-1}(t)$ is formally (since this is not point-wise convergent) our generalized Gaussian process: compare with $G^{1/2}(f)$ in (2.2) (up to a multiplicative constant as the control measures differ by a multiplicative constant $2$).
Processes with similar representations often arise as scaling limits of counting processes associated to models with certain renewal structures. See Alsmeyer et al., [1] for a limit theorem on the counting process for a family of decoupled renewal processes.